### v1

Abstract
Polychoric correlation is often an important building block in the analysis of rating data, particularly for structural equation models. However, the commonly employed maximum likelihood (ML) estimator is highly susceptible to misspecification of the polychoric correlation model, for instance through violations of latent normality assumptions. We propose a novel estimator that is designed to be robust to partial misspecification of the polychoric model, that is, the model is only misspecified for an unknown fraction of observations, for instance (but not limited to) careless respondents. In contrast to existing literature, our estimator makes no assumption on the type or degree of model misspecification. It furthermore generalizes ML estimation and is consistent as well as asymptotically normally distributed. We demonstrate the robustness and practical usefulness of our estimator in simulation studies and an empirical application on a Big Five administration. In the latter, the polychoric correlation estimates of our estimator and ML differ substantially, which, after further inspection, is likely due to the presence of careless respondents that the estimator helps identify.
Keywords: Polychoric correlation, model misspecification, robust estimation, careless responding, rating scales
1 Introduction
Ordinal data are ubiquitous in psychology and related fields. With such data, e.g., arising from responses to rating scales, it is often recommended to estimate correlation matrices through polychoric correlation coefficients (e.g., Foldnes & Grønneberg,, 2022; Garrido et al.,, 2013; Holgado-Tello et al.,, 2010). The resulting polychoric correlation matrix is an important building block in subsequent multivariate models like factor analysis models and structural equation models (SEMs), as well as in explanatory methods like principal component analysis, multidimensional scaling, and clustering techniques (see, e.g., Mair,, 2018, for an overview). An individual polychoric correlation coefficient is the population correlation between two underlying latent variables that are postulated to have generated the observed categorical data through an unobserved discretization process. Traditionally, it is assumed that the two latent variables are standard bivariate normally distributed (Pearson,, 1901) to estimate the polychoric correlation coefficient from observed ordinal data. Estimation of this latent normality model, called the polychoric model, is commonly carried out via maximum likelihood (Olsson,, 1979). However, recent work has demonstrated that maximum likelihood (ML) estimation of polychoric correlation is highly sensitive to violations of the assumption of underlying normality. Violations of this assumption result in a misspecified polychoric model, which can lead to substantially biased estimates of its parameters and those of subsequent multivariate models, particularly SEMs (Grønneberg & Foldnes,, 2022; Foldnes & Grønneberg,, 2022, 2020, 2019).
Motivated by the recent interest in non-robustness of ML, we study estimation of the polychoric model under a misspecification framework stemming from the robust statistics literature (e.g., Huber & Ronchetti,, 2009). In this setup, which we call partial misspecification here, the polychoric model is potentially misspecified for an unknown (and possibly zero-valued) fraction of observations. Heuristically, the model is misspecified such that the affected subset of observations contains little to no relevant information for the parameter of interest, the polychoric correlation coefficient. Examples of such uninformative observations include careless responses, misresponses, or responses due to item misunderstanding. Especially careless responding has been identified as a major threat to the validity of questionnaire-based research findings (e.g., Huang et al., 2015b, ; Meade & Craig,, 2012; Credé,, 2010; Woods,, 2006). We demonstrate that already a small fraction of uninformative observations (such as careless respondents) can result in considerably biased ML estimates.
As a remedy and our main contribution, we propose a novel way to estimate the polychoric model which is robust to partial model misspecification. Our estimator generalizes the ML estimator, but, in contrast to ML, does not crucially rely on correct specification of the model. Specifically, our estimator allows the model to be misspecified for an unknown fraction of uninformative responses in a sample, but makes no assumption whatsoever on the type, magnitude, or location of potential misspecification. The estimator is designed to identify such responses and to simultaneously downweigh their influence so that the polychoric model can be accurately estimated from the remaining responses generated by latent normality. Conversely, if the polychoric model is correctly specified, that is, latent normality holds true for all observations, our estimator and ML estimation are asymptotically equivalent. As such, our proposed estimator can be thought of as a generalized ML estimator that is robust to potential partial model misspecification, due to, for instance (but not limited to) careless responding. We show that our robust estimator is consistent and asymptotically normal under the polychoric model while possessing similar asymptotic properties under misspecification, and it comes at no additional computational cost compared to ML.
The partial misspecification framework in this paper is fundamentally different to that considered in recent literature on misspecified polychoric models. In this literature (e.g., Grønneberg & Foldnes,, 2022; Foldnes & Grønneberg,, 2022, 2020, 2019), the polychoric model is misspecified in the sense that all (unobserved) realizations of the latent continuous variables come from a distribution that is nonnormal. Under this framework, which is also known as distributional misspecification, the parameter of interest is the correlation coefficient of the latent nonnormal distribution, and all observations are informative for this parameter. While the distributional misspecification framework led to novel insights regarding (the lack of) robustness in ML estimation of polychoric correlation, the partial misspecification framework of this paper can provide complimentary insights regarding the effects of a fraction of uninformative observations in a sample (such as careless responses), which is our primary objective.
Nevertheless, while our estimator is designed to be robust to partial misspecification caused by some uninformative responses, it can in some situations also provide a robustness gain under distributional misspecification. It turns out that if a nonnormal latent distribution differs from a normal distribution mostly in the tails, our estimator produces less biased estimates than ML because it can downweigh observations that are father from the center.
To enhance accessibility and adoption by empirical researchers, an implementation of our proposed methodology in R (R Core Team,, 2024) is freely available in the package robcat (for “ROBust CATegorical data analysis”) at https://github.com/mwelz/robcat. [Replication files have been submitted for review together with this manuscript and will be made publicly available upon acceptance, with a link included here.]
This paper is structured as follows. We start with reviewing related literature (Section 2) followed by the polychoric correlation model and ML estimation thereof (Section 3). Afterwards, we elaborate on the partial misspecification framework (Section 4) and introduce our robust generalized ML estimator including its statistical properties (Section 5). These properties are then examined by a simulation study in which we vary the misspecification fraction systematically, and compare the result to the commonly employed standard ML estimator (Section 6). Subsequently, we demonstrate the practical usefulness in an empirical application on a Big Five administration (Goldberg,, 1992) by Arias et al., (2020), where we find evidence of careless responding, manifesting in differences in polychoric correlation estimates of as much as 0.3 between our robust estimator and ML (Section 7). We then investigate the performance of the estimator under distributional misspecification (Section 8) and conclude with a discussion of the results and avenues for further research (Section 9).
2 Related literature
In this section we briefly review the literature on misspecification of the polychoric model due to nonnormality of the underlying latent variables, as well as the literature on careless responding.
ML estimation of polychoric correlations was originally believed to be fairly robust to slight to moderate distributional misspecification (Li,, 2016; Coenders et al.,, 1997; Flora & Curran,, 2004; Maydeu-Olivares,, 2006). This belief was based on simulations that generated data for nonnormal latent variables via the Vale-Maurelli method (Vale & Maurelli,, 1983) that were then discretized to ordinal data. However, Grønneberg & Foldnes, (2019) show that the distribution of ordinal data generated in this way is indistinguishable from that of ordinal data stemming from discretizing normally distributed latent variables. In other words, simulation studies that ostensibly modeled nonnormality did in fact model normality. Simulating ordinal data in a way that ensures proper violations of latent normality (Grønneberg & Foldnes,, 2017) reveals that polychoric correlation is in fact highly susceptible to distributional misspecification, resulting in possibly large biases (Foldnes & Grønneberg,, 2020, 2022; Grønneberg & Foldnes,, 2022). Consequently, it is recommended to test for the validity of the latent normality assumption, for instance by using the bootstrap test of Foldnes & Grønneberg, (2020).
Another source of model misspecification occurs when the polychoric model is only misspecified for an uninformative subset of a sample (partial misspecification), where, in the context of this paper, the term “uninformative” refers to an absence of relevant information for polychoric correlation, for instance in careless responses. Careless responding “occurs when participants are not basing their response on the item content”, for instance when a participant is “unmotivated to think about what the item is asking” (Ward & Meade,, 2023). It has been shown to be a major threat to the validity of research results through a variety of psychometric issues, such as reduced scale reliability (Arias et al.,, 2020) and construct validity (Kam & Meyer,, 2015), attenuated factor loadings, improper factor structure, and deteriorated model fit in factor analyses (Arias et al.,, 2020; Huang et al., 2015a, ; Woods,, 2006), as well as inflated type I or type II errors in hypothesis testing (Arias et al.,, 2020; Huang et al., 2015b, ; Maniaci & Rogge,, 2014; McGrath et al.,, 2010; Woods,, 2006). Careless responding is widely prevalent (Ward & Meade,, 2023; Bowling et al.,, 2016; Meade & Craig,, 2012) with most estimates on its prevalence ranging from 10–15% of study participants (Curran,, 2016; Huang et al., 2015b, ; Huang et al.,, 2012; Meade & Craig,, 2012), while already a prevalence 5–10% can jeopardize the validity of research findings (Arias et al.,, 2020; Credé,, 2010; Woods,, 2006). In fact, Ward & Meade, (2023) conjecture that careless responding is likely present in all survey data. However, to the best of our knowledge, the effects of careless responding on estimates of the polychoric model have not yet been studied.
3 Polychoric correlation
The polychoric correlation model (Pearson & Pearson,, 1922) models the association between two discrete ordinal variables by assuming that an observed pair of responses to two polytomous items is governed by an unobserved discretization process of latent variables that jointly follow a bivariate standard normal distribution. If both items are dichotomous, the polychoric correlation model reduces to the tetrachoric correlation model of Pearson, (1901). In the following, we first define the model and review maximum likelihood (ML) estimation thereof and then introduce a robust estimator in the next section.
3.1 The polychoric model
For the ease of exposition, we restrict our presentation to the bivariate polychoric model. The model naturally generalizes to higher dimensions, see, e.g., Muthén, (1984).
Let there be two ordinal random variables, $X$ and $Y$, that take values in the sets $\mathcal{X}=\{1,2,\dots,K_{X}\}$ and $\mathcal{Y}=\{1,2,\dots,K_{Y}\}$, respectively. The assumption that the sets contain adjacent integers is without loss of generality. Suppose there exist two continuous latent random variables, $\xi$ and $\eta$, that govern the ordinal variables through the discretization model
$X=\begin{cases}1&\text{if }\xi<a_{1},\\ 2&\text{if }a_{1}\leq\xi<a_{2},\\ 3&\text{if }a_{2}\leq\xi<a_{3},\\ \vdots\\ K_{X}&\text{if }a_{K_{X}-1}\leq\xi,\\ \end{cases}$ and $Y=\begin{cases}1&\text{if }\eta<b_{1},\\ 2&\text{if }b_{1}\leq\eta<b_{2},\\ 3&\text{if }b_{2}\leq\eta<b_{3},\\ \vdots\\ K_{Y}&\text{if }b_{K_{Y}-1}\leq\eta,\\ \end{cases}$ (1)
where the fixed but unobserved parameters $a_{1}<a_{2}<\dots<a_{K_{X}-1}$ and $b_{1}<b_{2}<\dots<b_{K_{Y}-1}$ are called thresholds.
The primary object of interest is the population correlation between the two latent variables. To identify this quantity from the ordinal variables $(X,Y)$, one assumes that the continuous latent variables follow a standard bivariate normal distribution with unobserved pairwise correlation coefficient $\rho\in(-1,1)$, that is,
$\begin{pmatrix}\xi\\ \eta\end{pmatrix}\sim\textnormal{N}_{2}\Bigg{(}\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}\Bigg{)}.$ (2)
Combining the discretization model (1) with the latent normality model (2) yields the polychoric model. In this model, one refers to the correlation parameter $\rho=\mathrm{Cor}\left[\xi,\ \eta\right]$ as the polychoric correlation coefficient of the ordinal $X$ and $Y$. The polychoric model is subject to $d=K_{X}+K_{Y}-1$ parameters, namely the polychoric correlation coefficient from the latent normality model (2) and the two sets of thresholds from the discretization model (1). These parameters are jointly collected in a $d$-dimensional vector
$\bm{\theta}=\left(\rho,a_{1},a_{2},\dots,a_{K_{X}-1},b_{1},b_{2},\dots,b_{K_{Y}-1}\right)^{\top}.$
Under the polychoric model, the probability of observing an ordinal response $(x,y)\in\mathcal{X}\times\mathcal{Y}$ at a parameter vector $\bm{\theta}$ is given by
$p_{xy}(\bm{\theta})=\mathbb{P}_{\bm{\theta}}\left[X=x,Y=y\right]=\int_{a_{x-1}}^{a_{x}}\int_{b_{y-1}}^{b_{y}}\phi_{2}\left(t,s;\rho\right)\textnormal{d}s\ \textnormal{d}t,$ (3)
where we use the conventions $a_{0}=b_{0}=-\infty,a_{K_{X}}=b_{K_{Y}}=+\infty$, and
$\phi_{2}\left(u,v;\rho\right)=\frac{1}{2\pi\sqrt{1-\rho^{2}}}\exp\left(-\frac{u^{2}-2\rho uv+v^{2}}{2(1-\rho^{2})}\right)$
denotes the density of the standard bivariate normal distribution function with correlation parameter $\rho\in(-1,1)$ at some $u,v\in\mathbb{R}$, with corresponding distribution function
$\Phi_{2}\left(u,v;\rho\right)=\int_{-\infty}^{u}\int_{-\infty}^{v}\phi_{2}\left(t,s;\rho\right)\textnormal{d}s\ \textnormal{d}t.$
To distinguish arbitrary parameters $\bm{\theta}$ from a specific value under which the polychoric model generates ordinal data, denote the latter by $\bm{\theta}_{*}=\left(\rho_{*},a_{*,1},\dots,a_{*,K_{X}-1},b_{*,1},\dots,b_{*,K_{Y}-1}\right)^{\top}$. Given a random sample of ordinal data generated by a polychoric model under parameter value $\bm{\theta}_{*}$, the statistical problem is to estimate the true $\bm{\theta}_{*}$, which is traditionally achieved by the maximum likelihood estimator of Olsson, (1979).
3.2 Maximum likelihood estimation
Suppose we observe a sample $\{(X_{i},Y_{i})\}_{i=1}^{N}$ of $N$ independent copies of $(X,Y)$ generated by the polychoric model under the true parameter $\bm{\theta}_{*}$. The sample may be observed directly or as a $K_{X}\times K_{Y}$ contingency table that cross-tabulates the observed frequencies. Denote by
$N_{xy}=\sum_{i=1}^{N}\mathds{1}\left\{X_{i}=x,Y_{i}=y\right\}$
the observed empirical frequency of a response $(x,y)\in\mathcal{X}\times\mathcal{Y}$, where the indicator function $\mathds{1}\left\{E\right\}$ takes value 1 if an event $E$ is true, and 0 otherwise. The maximum likelihood estimator (MLE) of $\bm{\theta}_{*}$ can be expressed as
$\widehat{\bm{\theta}}_{N}^{\mathrm{\ MLE}}=\arg\max_{\bm{\theta}\in\bm{\Theta}}\bigg{\{}\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}N_{xy}\log\left(p_{xy}(\bm{\theta})\right)\bigg{\}},$ (4)
where the $p_{xy}(\bm{\theta})$ are the response probabilities in (3), and
$\bm{\Theta}=\bigg{(}\Big{(}\rho,\big{(}a_{i}\big{)}_{i=1}^{K_{X}-1},\big{(}b_{j}\big{)}_{j=1}^{K_{Y}-1}\Big{)}^{\top}\ \Big{|}\ \rho\in(-1,1),\ a_{1}<\dots<a_{K_{X}-1},\ b_{1}<\dots<b_{K_{Y}-1}\bigg{)}$ (5)
is a set of parameters $\bm{\theta}$ that rules out degenerate cases such as $\rho=\pm 1$ or thresholds that are not strictly monotonically increasing. This estimator, its computational details, as well as its statistical properties are derived in Olsson, (1979). In essence, if the polychoric model (1) is correctly specified—that is, the underlying latent variables $(\xi,\eta)$ are indeed standard bivariate normal—then the estimator $\widehat{\bm{\theta}}_{N}^{\mathrm{\ MLE}}$ is consistent for the true $\bm{\theta}_{*}$. In addition, $\widehat{\bm{\theta}}_{N}^{\mathrm{\ MLE}}$ is asymptotically normally distributed with mean zero and covariance matrix being equal to the model’s inverse Fisher information matrix, which makes it fully efficient.
As a computationally attractive alternative to estimating all parameters in $\bm{\theta}_{*}$ simultaneously in problem (4), one may consider a “2-step-approach” where only the correlation coefficient $\rho_{*}$ is estimated via ML, but not the thresholds. In this approach, one estimates in a first step the thresholds as quantiles of the univariate standard normal distribution, evaluated at the observed cumulative marginal proportion of each contingency table cell. Formally, in the 2-step-approach, thresholds $a_{*,x}$ and $b_{*,y}$ are respectively estimated via
$\widehat{a}_{x}=\Phi_{1}^{-1}\left(\frac{1}{N}\sum_{k=1}^{x}\sum_{y\in\mathcal{Y}}N_{ky}\right)\qquad\text{and}\qquad\widehat{b}_{y}=\Phi_{1}^{-1}\left(\frac{1}{N}\sum_{x\in\mathcal{X}}\sum_{l=1}^{y}N_{xl}\right),$ (6)
for $x=1,\dots,K_{X}-1$ and $y=1,\dots,K_{Y}-1$, where $\Phi_{1}^{-1}(\cdot)$ denotes the quantile function of the univariate standard normal distribution. Then, taking these threshold estimates as fixed in the polychoric model, one estimates in a second step the only remaining parameter, correlation coefficient $\rho_{*}$, via ML. The main advantage of the 2-step approach is reduced computational time, while it comes at the cost of being theoretically non-optimal because ML standard errors do not apply to the threshold estimators in (6) (Olsson,, 1979).
4 Conceptualizing model misspecification
To study the effects of partial model misspecification from a theoretical perspective, we first rigorously define this concept. In addition, we explain how it is different from distributional misspecification. This section is built on classic work on conceptualizing model misspecification in robust statistics (e.g., Huber & Ronchetti,, 2009; Hampel et al.,, 1986; Huber,, 1964).
4.1 Partial misspecification of the polychoric model
The polychoric model is partially misspecified when not all unobserved realizations of the latent variables $(\xi,\eta)$ come from a standard bivariate normal distribution. Specifically, we consider a situation where only a fraction $(1-\varepsilon)$ of those realizations are generated by a standard bivariate normal distribution with true correlation parameter $\rho_{*}$, whereas a fixed fraction $\varepsilon\in[0,0.5)$ are generated by some different but unspecified distribution $H$. Note that $H$ being unspecified allows its correlation coefficient to differ from $\rho_{*}$ so that realizations generated by $H$ may be uninformative for the true polychoric correlation coefficient $\rho_{*}$, such as, after discretization, careless responses, misresponses or responses stemming from item misunderstanding.
Formally, we say that the polychoric model is partially misspecified if the latent variables $(\xi,\eta)$ are jointly distributed according to the mixture distribution
$(u,v)\mapsto G_{\varepsilon}(u,v)=(1-\varepsilon)\Phi_{2}\left(u,v;\rho_{*}\right)+\varepsilon H(u,v),$ (7)
[…]
Leaving misspecifying distribution $H$ and misspecification fraction $\varepsilon$ unspecified in mixture distribution (7) means that we are not making any assumptions on the degree, magnitude, or type of misspecification (which is possibly absent altogether). Hence, in our context of responses to rating items, the polychoric model can be misspecfied due to an unlimited variety of reasons, for instance but not limited to careless/inattentive responding (e.g., straightlining, pattern responding, random-like responding), misresponses, or item misunderstanding.
The maximum misspecification fraction $\varepsilon$ in mixture distribution (7) is restricted to be less than 0.5 such that the polychoric model is correctly specified for the majority of observations, which is standard in the robust statistics literature (e.g., Hampel et al.,, 1986, p. 67). While it is in principle possible to also consider a misspecification fraction between $0.5$ and $1$, one would need to impose certain additional assumptions on the correct model to distinguish it from incorrect ones when the majority of observations are not generated by the correct model. Since we prefer refraining from imposing additional assumptions, we only consider $\varepsilon\in[0,0.5)$.
Furthermore, as another, more practical reason for considering $\varepsilon\in[0,0.5)$, having more than half of all observations in a sample being not informative for the quantity of interest would be indicative of serious data quality issues. When data quality is unreasonably low, it is doubtful whether the data are suitable for modeling analyses in the first place.
4.2 Response probabilities under partial misspecification
Under misspecified distribution $G_{\varepsilon}$ with misspecification fraction $\varepsilon\in[0,0.5)$, the probability of observing an ordinal response $(x,y)\in\mathcal{X}\times\mathcal{Y}$ is given by
$f_{\varepsilon}\left(x,y\right)=\mathbb{P}_{G_{\varepsilon}}\left[X=x,Y=y\right]=(1-\varepsilon)p_{xy}(\bm{\theta}_{*})+\varepsilon\int_{a_{\varepsilon,x-1}}^{a_{\varepsilon,x}}\int_{b_{\varepsilon,y-1}}^{b_{\varepsilon,y}}\textnormal{d}H,$ (8)
where the unobserved thresholds $a_{\varepsilon,x},b_{\varepsilon,y}$ discretize the fraction $\varepsilon$ of latent variables for which the polychoric model is misspecified. The thresholds $a_{\varepsilon,x},b_{\varepsilon,y}$ may be different from the true $a_{*,x},b_{*,y}$ and/or depend on misspecification fraction $\varepsilon$. However, it turns out that from a theoretical perspective, studying the case where the $a_{\varepsilon,x},b_{\varepsilon,y}$ are different from the $a_{*,x},b_{*,y}$ is equivalent to a case where they are equal. [^2] Let $H^{\prime}$ be an arbitrary misspecifying distribution of the latent variables. Let the thresholds that discretize these latent variables be given by arbitrary values $a^{\prime}_{\varepsilon,x},b^{\prime}_{\varepsilon,y}$, for $x\in\mathcal{X},y\in\mathcal{Y}$. Since one makes no assumption on the misspecifying distribution in (7), we can find another misspecifying distribution $H\neq H^{\prime}$ that, when discretized with the true thresholds $a_{*,x},b_{*,y}$, yields the same discretization as $H^{\prime}$ with thresholds $a^{\prime}_{\varepsilon,x},b^{\prime}_{\varepsilon,y}$. Formally, $\forall a^{\prime}_{\varepsilon,x},b^{\prime}_{\varepsilon,y},H^{\prime}\ \exists H\text{ s.t. }\int_{a^{\prime}_{\varepsilon,x-1}}^{a^{\prime}_{\varepsilon,x}}\int_{b^{\prime}_{\varepsilon,y-1}}^{b^{\prime}_{\varepsilon,y}}\textnormal{d}H^{\prime}=\int_{a_{*,x-1}}^{a_{*,x}}\int_{b_{*,y-1}}^{b_{*,y}}\textnormal{d}H$.
The population response probabilities $f_{\varepsilon}\left(x,y\right)$ in (8) are unknown in practice because they depend on unspecified and unmodeled quantities, namely the misspecifying distribution $H$, misspecification fraction $\varepsilon$, and the discretization thresholds of the latter.
Figure 1: Simulated data with $K_{X}=K_{Y}=5$ response options where the polychoric model is misspecified due to a fraction $\varepsilon=0.15$ of nonnormal data. The gray dots represent random draws of $(\xi,\eta)$ from the polychoric model with $\rho_{*}=0.5$, whereas the orange dots represent draws from a different distribution that inflates contingency table cell $(x,y)=(5,1)$. The blue lines indicate the location of the thresholds. In each cell, the numbers in parentheses denote the population probability of that cell under the true polychoric model.
Figure 1 visualizes a simulated example of bivariate data drawn from misspecified distribution $G_{\varepsilon}$, where a fraction of $\varepsilon=0.15$ of the data follow a (misspecifying) distribution with mean $(2,-2)$ (orange dots), whereas the remaining data are generated by a standard bivariate normal distribution with correlation $\rho_{*}=0.5$ (gray dots). In this example, the data from the misspecifying distribution will primarily inflate the contingency table cell $(x,y)=(5,1)$, in the sense that this cell will have a larger empirical frequency than the polychoric model allows for, since the probability of this cell is nearly zero at the polychoric model yet many realized responses will populate it. Consequently, due to misspecification of the (polychoric) model, a maximum likelihood estimate of $\rho_{*}$ on these data might be substantially biased for $\rho_{*}$. Indeed, calculating the ML estimator using the data plotted in Figure 1 yields an estimate of $\widehat{\rho}_{N}^{\mathrm{\ MLE}}=-0.05$, which is far off from the true $\rho_{*}=0.5$. In contrast, our proposed robust estimator, which is calculated from the exact same information as the ML estimator and is defined in the next section (Section 5), yields a fairly accurate estimate of $0.45$.
It is worth addressing that there exist nonnormal distributions of the latent variables $(\xi,\eta)$ that, after discretization with the same thresholds, result in the same response probabilities as under latent normality (Foldnes & Grønneberg,, 2019). This implies that there may exist misspecifying distributions $H$ and (nonzero) misspecification fractions $\varepsilon$ under which the population probabilities $f_{\varepsilon}\left(x,y\right)$ in (8) are equal to the true population probabilities of the polychoric model, $p_{xy}(\bm{\theta}_{*})$, that is, $f_{\varepsilon}\left(x,y\right)=p_{xy}(\bm{\theta}_{*})$ for $\varepsilon>0$. In this situation, the polychoric model is misspecified, but the misspecification does not have consequences because the response probabilities remain unaffected so that there is no robustness issue with ML estimation. To avoid cumbersome notation in the theoretical analysis of our robust estimator, we assume consequential misspecification throughout this paper, that is, $f_{\varepsilon}\left(x,y\right)\neq p_{xy}(\bm{\theta}_{*})$ whenever $\varepsilon>0$. However, it is silently understood that misspecification need not be consequential, in case of which there is no non-robustness issue and both the ML estimator and our robust estimator will be consistent for the true $\bm{\theta}_{*}$.
4.3 Distributional misspecification
A model is distributionally misspecified when all observations in a given sample are generated by a distribution that is different from the model distribution. In the context of the polychoric model, this means that all ordinal observations are generated by a latent distribution that is nonnormal. Let $G$ denote the unknown nonnormal distribution that the latent variables $(\xi,\eta)$ jointly follow under distributional misspecification. The object of interest is the population correlation between latent $\xi$ and $\eta$ under distribution $G$, for which the normality-based ML estimator of Olsson, (1979) turns out to be substantially biased in many cases (e.g., Foldnes & Grønneberg,, 2020, 2022). As such, distributional misspecification is fundamentally different from partial misspecification: In the former, one attempts to estimate the population correlation of the nonnormal and unknown distribution that generated a sample, instead of estimating the polychoric correlation coefficient (which is the correlation under standard bivariate normality). In the latter, one attempts to estimate the polychoric correlation coefficient with a sample that has only been partly generated by latent normality (that is, the polychoric model). The assumption that the polychoric model is only partially misspecified for some uninformative observations enables one to still estimate the polychoric correlation coefficient of that model, which would not be feasible under distributional misspecification (at least not without additional assumptions).
Despite not being designed for distributional misspecification, the robust estimator introduced in the next section (Section 5) can offer a robustness gain in some situations where the polychoric model is distributionally misspecified. We discuss this in more detail in Section 8.
5 Robust estimation of polychoric correlation
The consistency of ML estimation of an arbitrary model crucially depends on correct specification of that model. Indeed, ML estimation can be severely biased even when the assumed model is only slightly misspecified (e.g., Huber & Ronchetti,, 2009; Hampel et al.,, 1986; Huber,, 1964). For instance, in many models of continuous variables like regression models, one single observation from a different distribution can be enough to make the ML estimator converge to an arbitrary value (Huber & Ronchetti,, 2009, see also Alfons et al.,, 2022, for the special case of mediation analysis). The non-robustness of ML estimation of the polychoric model has been demonstrated empirically by Foldnes & Grønneberg, (2020, 2022); Grønneberg & Foldnes, (2022) for the case of distributional misspecification. In this section, we introduce an estimator that is designed to be robust to partial misspecification when present, but remains (asymptotically) equivalent to the ML estimator of Olsson, (1979) when misspecification is absent. We furthermore derive the statistical properties of the proposed estimator.
Throughout this section, let $\{(X_{i},Y_{i})\}_{i=1}^{N}$ be an observed $N$-sized ordinal sample generated by discretizing latent variables $(\xi,\eta)$ that follow the unknown and unspecified mixture distribution $G_{\varepsilon}$ in (7). Hence, the polychoric model is possibly misspecified for an unknown fraction $\varepsilon$ of the sample.
5.1 The estimator
The proposed estimator is a special case of a robust estimator for general models of categorical data introduced by Welz, (2024), and is based on the following idea. The empirical relative frequency of a response $(x,y)\in\mathcal{X}\times\mathcal{Y}$, denoted
$\widehat{f}_{N}(x,y)=N_{xy}/N=\frac{1}{N}\sum_{i=1}^{N}\mathds{1}\left\{X_{i}=x,Y_{i}=y\right\},$
is a consistent nonparametric estimator of the population response probability in (8),
$f_{\varepsilon}\left(x,y\right)=\mathbb{P}_{G_{\varepsilon}}\left[X=x,Y=y\right]=(1-\varepsilon)p_{xy}(\bm{\theta}_{*})+\varepsilon\int_{a_{\varepsilon,x-1}}^{a_{\varepsilon,x}}\int_{b_{\varepsilon,y-1}}^{b_{\varepsilon,y}}\textnormal{d}H,$
as $N\to\infty$ (see, e.g., Chapter 19.2 in Van der Vaart,, 1998). If the polychoric model is correctly specified $(\varepsilon=0)$, then $\widehat{f}_{N}(x,y)$ will converge to the true model probability $p_{xy}(\bm{\theta}_{*})$ because
$f_{0}(x,y)=p_{xy}(\bm{\theta}_{*}).$
Conversely, if the polychoric model is misspecified $(\varepsilon>0)$, then $\widehat{f}_{N}(x,y)$ does not converge to the true $p_{xy}(\bm{\theta}_{*})$ because
$f_{\varepsilon}\left(x,y\right)\neq p_{xy}(\bm{\theta}_{*}),$
since we assume consequential misspecification.
It follows that if the polychoric model is misspecified, there exists no parameter value $\bm{\theta}\in\bm{\Theta}$ whose associated model probability $p_{xy}(\bm{\theta})$ the nonparametric estimate $\widehat{f}_{N}(x,y)$ converges to. Hence, it is indicative of model misspecification if $\widehat{f}_{N}(x,y)$ does not converge to any polychoric model probability $p_{xy}(\bm{\theta})$, resulting in a discrepancy of $\widehat{f}_{N}(x,y)$ and $p_{xy}(\bm{\theta})$. This observation can be exploited in model fitting by minimizing the discrepancy between the empirical relative frequencies, $\widehat{f}_{N}(x,y)$, and theoretical model probabilities, $p_{xy}(\bm{\theta})$, to find the most accurate fit that can be achieved with the polychoric model for the observed data. Specifically, our estimator minimizes with respect to $\bm{\theta}$ the loss function
$L\left(\bm{\theta},\ \widehat{f}_{N}\right)=\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}\varphi\left(\frac{\widehat{f}_{N}(x,y)}{p_{xy}(\bm{\theta})}\right)p_{xy}(\bm{\theta}),$ (9)
where $\varphi:[0,\infty)\to\mathbb{R}$ is a prespecified discrepancy function that will be defined momentarily. The proposed estimator $\widehat{\bm{\theta}}_{N}$ is given by the value minimizing the objective loss over parameter space $\bm{\Theta}$,
$\widehat{\bm{\theta}}_{N}=\arg\min_{\bm{\theta}\in\bm{\Theta}}L\Big{(}\bm{\theta},\widehat{f}_{N}\Big{)}.$ (10)
For the discrepancy function choice $\varphi(z)=\varphi^{\mathrm{MLE}}(z)=z\log(z)$, it can be easily verified that $\widehat{\bm{\theta}}_{N}$ coincides with the MLE $\widehat{\bm{\theta}}_{N}^{\mathrm{\ MLE}}$ in (4). In the following, we motivate a specific choice of discrepancy function $\varphi(\cdot)$ that makes the estimator $\widehat{\bm{\theta}}_{N}$ less susceptible to misspecification of the polychoric model while preserving equivalence with ML estimation in the absence of misspecification.
The centralized fraction between empirical relative frequencies and model probabilities, $\widehat{f}_{N}(x,y)/p_{xy}(\bm{\theta})-1$, is referred to as Pearson residual (Lindsay,, 1994) and can be interpreted as a goodness-of-fit measure. In our notation, it is simpler to renounce on subtracting the value 1 so that we call
$\frac{\widehat{f}_{N}(x,y)}{p_{xy}(\bm{\theta})}$
a decentralized Pearson residual (DPR), which takes values in $[0,+\infty)$. DPR values close to 1 indicate a good fit between data and polychoric model at $\bm{\theta}$, whereas values toward 0 or +$\infty$+ ∞ indicate a poor fit because empirical response probabilities disagree with their model counterparts. To achieve robustness to misspecification of the polychoric model, responses that cannot be modeled well by the polychoric model, as indicated by their DPR being away from 1, should receive less weight in the estimation procedure such that they do not over-proportionally affect the fit. Downweighting when necessary is achieved through a specific choice of discrepancy function $\varphi(\cdot)$ proposed by Welz, (2024), which is a special case of a function suggested by Ruckstuhl & Welsh, (2001). The discrepancy function reads
$\varphi(z)=\begin{cases}z\log(z)&\textnormal{ if }z\in[0,c],\\ z(\log(c)+1)&\textnormal{ if }z>c,\end{cases}$ (11)
where $c\in[1,\infty]$ is a prespecified tuning constant that governs the estimator’s behavior at the DPR of each possible response. Figure 2 visualizes this function for the example choice $c=1.6$ as well as the ML discrepancy function function $\varphi^{\mathrm{MLE}}(z)=z\log(z)$.
Figure 2: Visualization of the robust discrepancy function $\varphi(z)$ in (11) for $c=1.6$ (solid line) and the ML discrepancy function $\varphi^{\mathrm{MLE}}(z)=z\log(z)$ (dotted line).
For the choice $c=+\infty$, minimizing the loss (9) is equivalent to maximizing the log-likelihood objective in (4), meaning that the estimator $\widehat{\bm{\theta}}_{N}$ is equal to $\widehat{\bm{\theta}}_{N}^{\mathrm{\ MLE}}$ for this choice of $c$. More specifically, if a DPR $z=\frac{\widehat{f}_{N}(x,y)}{p_{xy}(\bm{\theta})}$ of a response $(x,y)\in\mathcal{X}\times\mathcal{Y}$ is such that $z\in[0,c]$ for fixed $c\geq 1$, then the estimation procedure behaves at this response like in classic ML estimation. As argued before, in the absence of misspecification, $\widehat{f}_{N}(x,y)$ converges to $p_{xy}(\bm{\theta}_{*})$ for all responses $(x,y)\in\mathcal{X}\times\mathcal{Y}$, therefore all DPRs are asymptotically equal to 1. Hence, if the polychoric model is correctly specified, then estimator $\widehat{\bm{\theta}}_{N}$ is asymptotically equivalent to the MLE $\widehat{\bm{\theta}}_{N}^{\mathrm{\ MLE}}$ for any tuning constant value $c\geq 1$. On the other hand, if a response’s DPR $z$ is larger than $c$, that is, $z>c\geq 1$, then the estimation procedure does not behave like in ML, but the response’s contribution to loss (9) is linear rather than super-linear like in ML (Figure 2). It follows that the influence of responses that cannot be fitted well by the polychoric model is downweighted to prevent them from dominating the fit. The tuning constant $c\geq 1$ is the “threshold” beyond which a DPR will be downweighted, so the choice thereof determines what is considered an insufficient fit. The closer to 1 the tuning constant $c$ is chosen, the more robust the estimator.
With the proposed choice of $\varphi(\cdot)$, we stress that our estimator $\widehat{\bm{\theta}}_{N}$ in (10) has the same time complexity as ML, that is, $O\big{(}K_{X}\cdot K_{Y}\big{)}$, since one needs to calculate the Pearson residual of all $K_{X}\cdot K_{Y}$ possible responses for every candidate parameter value. Consequently, our proposed estimator does not incur any additional computational cost compared to ML.
5.2 Statistical properties
We first address what the generalized ML estimator $\widehat{\bm{\theta}}_{N}$ in (10) actually estimates. Its estimand is given by
$\bm{\theta}_{0}=\arg\min_{\bm{\theta}\in\bm{\Theta}}L\Big{(}\bm{\theta},f_{\varepsilon}\Big{)}.$
This minimization problem is simply the population analogue to the minimization problem in (10) that the sample-based $\widehat{\bm{\theta}}_{N}$ solves because the probabilities $f_{\varepsilon}\left(x,y\right)$ are the population analogues to the empirical probabilities $\widehat{f}_{N}(x,y)$.
If the polychoric model is correctly specified, the estimand $\bm{\theta}_{0}$ equals the true parameter $\bm{\theta}_{*}$. Indeed, if $\varepsilon=0$, then $f_{0}(x,y)=p_{xy}(\bm{\theta}_{*})$ for all $(x,y)\in\mathcal{X}\times\mathcal{Y}$, so it follows that the loss
$\begin{split}&L\Big{(}\bm{\theta},f_{0}\Big{)}\\ &\quad=\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}\varphi\left(\frac{p_{xy}(\bm{\theta}_{*})}{p_{xy}(\bm{\theta})}\right)p_{xy}(\bm{\theta})\\ &\quad=\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}p_{xy}(\bm{\theta}_{*})\Bigg{[}{\log}\Bigg{(}\frac{p_{xy}(\bm{\theta}_{*})}{p_{xy}(\bm{\theta})}\Bigg{)}\mathds{1}\left\{\frac{p_{xy}(\bm{\theta}_{*})}{p_{xy}(\bm{\theta})}\in[0,c]\right\}+\left(\log(c)+1\right)\mathds{1}\left\{\frac{p_{xy}(\bm{\theta}_{*})}{p_{xy}(\bm{\theta})}>c\right\}\Bigg{]}\end{split}$
attains its global minimum of zero if and only if $\bm{\theta}=\bm{\theta}_{*}$, for any choice of $c\geq 1$. Thus, in the absence of misspecification, our estimator estimates the same quantity as the MLE, namely the true $\bm{\theta}_{*}$.
Figure 3: The population estimand $\rho_{0}$ of the polychoric correlation coefficient for various degrees of misspecification fractions $\varepsilon$ ($x$-axis) and tuning constants $c$ (line colors). The ML estimand corresponds to $c=+\infty$. There are $K_{X}=K_{Y}=5$ response options and the true value corresponds to $\rho_{*}=0.5$ (dashed line). The misspecification type here is the same as in Figure 1.
However, in the presence of misspecification ($\varepsilon>0$), the estimand $\bm{\theta}_{0}$ is generally different from the true $\bm{\theta}_{*}$. The population estimand being different from the true value translates to biased estimates of the latter as a consequence of the misspecification. How much they are apart depends on the unknown fraction of misspecification $\varepsilon$, the unknown type of misspecification $H$, as well as the choice of tuning constant $c$ in $\varphi(\cdot)$. Mainly, the larger $\varepsilon$ (more severe misspecification) and $c$ (less downweighting of hard-to-fit responses), the further $\bm{\theta}_{0}$ is away from $\bm{\theta}_{*}$. Figure 3 illustrates this behavior for the polychoric correlation coefficient at an example misspecified distribution that is described in Section 4.1 and in which the true polychoric correlation under the correct model amounts to  $\rho_{*}=0.5$. For increasing misspecification fractions, the MLE ($c=+\infty)$ estimates a parameter value that is increasingly farther away from the true $\bm{\theta}_{*}$, where already a misspecification fraction of less than $\varepsilon=0.15$ suffices for a sign flip in the correlation coefficient. Conversely, choosing tuning constant $c$ to be near 1 results in a much less severe bias. For instance, even at misspecification degree $\varepsilon=0.3$, the difference between estimand and true value is less than approximately 0.1.
Overall, finite choices of $c$ lead to an estimator that is at least as accurate as the MLE, and more accurate under misspecification of the polychoric model, thereby gaining robustness to misspecification.
A relevant question is whether the true parameter $\bm{\theta}_{*}$ can be recovered when $\varepsilon>0$ such that it can be estimated using $\widehat{\bm{\theta}}_{N}$ combined with a bias correction term. To derive such a bias correction term, one would need to impose assumptions on the misspecification fraction and type of misspecification. However, if one has strong prior beliefs about how the polychoric model is misspecified, modeling them explicitly rather than relying on the polychoric model seems more appropriate. Yet, one’s beliefs about misspecification may not be accurate, so attempts to explicitly model the misspecification may themselves result in a misspecified model. Consequently, robust estimation traditionally refrains from making assumptions on how a model may potentially be misspecified by leaving $\varepsilon$ and $H$ unspecified in mixture distribution (7). Instead, one may use a robust estimator to identify data points that cannot be modeled with the model at hand, like the one presented in this paper.
It can be shown that under certain regularity conditions that do not restrict the degree or type of possible partial misspecification beyond $\varepsilon\in[0,0.5)$, the robust estimator $\widehat{\bm{\theta}}_{N}$ is consistent for estimand $\bm{\theta}_{0}$ as well as asymptotically normally distributed. With this limit theory, one can construct standard errors and confidence intervals for every element in $\bm{\theta}_{0}$. Importantly, in the absence of model misspecification (such that $\bm{\theta}_{0}=\bm{\theta}_{*}$), the asymptotic covariance matrix of the robust estimator is equal to that of the fully efficient ML estimator. Hence, if the model is correctly specified, then robust estimator and ML estimator are asymptotically first and second order equivalent. We refer to Appendix A for a rigorous exposition and discussion of the robust estimator’s asymptotic properties.
5.3 Implementation
We provide a free and open source implementation of our proposed methodology in a package for the statistical programming environment R (R Core Team,, 2024). The package is called robcat (for “ROBust CATegorical data analysis”) and is available at https://github.com/mwelz/robcat. To maximize speed and performance, the package is predominantly developed in C++ and integrated to R via Rcpp (Eddelbuettel,, 2013). All numerical results in this paper were obtained with this package.
The estimator’s minimization problem in (10) can be solved with standard algorithms for numerical optimization. Yet, for the robust estimator we recommend to use constrained optimization techniques for imposing the boundary constraint on the correlation coefficient and the monotonicity constraints on the thresholds; see (5). In package robcat, the simplex algorithm of Nelder & Mead, (1965) for constrained optimization is used for this purpose.
An important user choice is that of the tuning constant $c$ in discrepancy function (11). The closer $c$ is to 1, the more robust the estimator will be to possible misspecification of the polychoric model (see, e.g., Figure 3). However, in the presence of model misspecification, the more robust the estimator is made, the larger its estimation variance becomes. This is an instance of a well-known tradeoff between robustness and efficiency (e.g., Huber & Ronchetti,, 2009). Here, the choice of $c$ governs the robustness-efficiency tradeoff. In simulation experiments, we found that the choice $c=1.6$ yielded a good compromise (see, e.g., Figure 3 for the robustness gain), so we choose to use this value for all applications in this paper. However, we acknowledge that a detailed study, possibly a theoretical one, is necessary to provide guidelines on the choice of $c$. We will explore this in future work.
Furthermore, a two-step estimation procedure like in (6) is not recommended for robust estimation. The possible presence of responses that have not been generated by the polychoric model can inflate the empirical cumulative marginal proportion of some responses, which may result in a sizable bias of threshold estimates (6), possibly translating into biased estimates of polychoric correlation coefficients in the second stage. Our robust estimator therefore estimates all model parameters (thresholds and polychoric correlation) simultaneously.
6 Simulation studies on partial misspecification
In this section, we employ two simulation studies to demonstrate the robustness gain of our proposed estimator under partial misspecification of the polychoric model. The first simulation design is intended to illustrate the performance of the robust estimator for the estimation of an individual polychoric correlation coefficient. The second design considers estimation of a polychoric correlation matrix, which is often am important building block in subsequent multivariate analyses, such a structural equation modeling or principal component analyses.
6.1 Individual polychoric correlation coefficient
Let there be $K_{X}=K_{Y}=5$ response categories for each of the two rating variables and define the true thresholds in the discretization process (1) as
$a_{*,1}=b_{*,1}=-1.5,\quad a_{*,2}=b_{*,2}=-0.5,\quad a_{*,3}=b_{*,3}=0.5,\quad a_{*,4}=b_{*,4}=1.5,$
and let the true polychoric correlation coefficient in latent normality model (2) be $\rho_{*}=0.5$. To simulate partial misspecification of the polychoric model, we let a fraction $\varepsilon$ of the data be generated by a misspecifying bivariate normal distribution with mean $(2,-2)^{\top}$, variances $(0.2,0.2)^{\top}$, and zero covariance (and therefore zero correlation). This misspecifying distribution will inflate the empirical frequency of contingency table cells $(x,y)\in\{(5,1),(4,3),(5,2)\}$, in the sense that they have a higher realization probability than under the true polychoric model. In fact, the data plotted in Figure 1 were generated by this process for misspecification fraction $\varepsilon=0.15$, and one can see in this figure that particularly cell $(x,y)=(5,1)$ is sampled quite frequently although it only has a near-zero probability at the true polychoric model. The data points causing these three cells to be inflated are instances of negative leverage points. Here, such leverage points drag correlational estimates away from a positive value towards zero or, if there are sufficiently many of them, even a negative value. One may think of such points as the responses of careless or inattentive participants whose responses are not based on item content.
For misspecification fraction $\varepsilon\in\{0,0.1,0.2\}$, we sample $N=1,000$ ordinal responses from this data generating process and estimate the true parameter $\bm{\theta}_{*}$ with the MLE (Olsson,, 1979), our proposed estimator with tuning constant set to $c=1.6$, and, for comparison, the Pearson sample correlation calculated on the integer-valued responses. We repeat this procedure for 1,000 simulated datasets. As performance measures, we calculate the average bias, standard deviation across repetitions, coverage, and length of confidence intervals at significance level $\alpha=0.05$. The coverage is defined as proportion of $(1-\alpha)$-th confidence intervals $[\widehat{\rho}_{N}\mp q_{1-\alpha/2}\cdot\text{SE}(\widehat{\rho}_{N})]$ that contain the true $\rho_{*}$, where $q_{1-\alpha/2}$ is the $(1-\alpha/2)$-th quantile of the standard normal distribution and $\text{SE}(\widehat{\rho}_{N})$ is the standard error of $\widehat{\rho}_{N}$, which is constructed using the limit theory developed in Theorem 1 in Appendix A. The length of a confidence interval is given by $2\cdot q_{1-\alpha/2}\cdot\text{SE}(\widehat{\rho}_{N})$.
Figure 4: Boxplot visualization of the bias of three estimators of the polychoric correlation coefficient, $\widehat{\rho}_{N}-\rho_{*}$, for various degrees of misspecification of the polychoric model across 1,000 replications. The estimators are the proposed robust estimator with $c=1.6$ (left), the MLE of Olsson, (1979) (center), and the Pearson sample correlation coefficient (right).
Misspecification Estimator $\widehat{\rho}_{N}$ Bias StDev Coverage CI length
$\varepsilon=0$ Polycor robust 0.504 0.004 0.027 0.930 0.104
Polycor ML 0.500 0.000 0.026 0.943 0.102
Sample cor 0.457 $-0.043$ 0.025 0.702 0.110
$\varepsilon=0.1$ Polycor robust 0.466 $-0.034$ 0.038 0.911 0.152
Polycor ML 0.097 $-0.403$ 0.029 0.000 0.134
Sample cor 0.084 $-0.416$ 0.026 0.000 0.124
$\varepsilon=0.2$ Polycor robust 0.439 $-0.061$ 0.051 0.951 0.220
Polycor ML $-0.172$ $-0.672$ 0.028 0.000 0.133
Sample cor $-0.133$ $-0.633$ 0.026 0.000 0.123
Table 1: Performance measures of of the three estimators, Pearson sample correlation, the MLE, and the proposed robust estimator with $c=1.6$, for various misspecification fractions across 1,000 simulated datasets. The true polychoric correlation coefficient is $\rho_{*}=0.5$. The performance measures are the average point estimate of the polychoric correlation coefficient, $\widehat{\rho}_{N}$, average bias ($\widehat{\rho}_{N}-\rho_{*}$), the standard deviation of the $\widehat{\rho}_{N}$ (“StDev”), the estimator’s coverage with respect to the true $\rho_{*}$ at significance level $\alpha=0.05$, as well as the length of the estimator’s confidence interval, again at level $\alpha=0.05$.
Figure 4 visualizes the bias of each estimator with respect to the true polychoric correlation $\rho_{*}$ across the  1,000 simulated datasets. An analogous plot for the whole parameter $\bm{\theta}_{*}$ can be found in Appendix C; the results are similar to those of $\rho_{*}$. For all considered misspecification fractions, the estimates of the MLE and sample correlation are similar, which is expected because these two estimators are known to yield similar results when there are five or more rating options and the discretization thresholds are symmetric (cf., Rhemtulla et al.,, 2012). In the absence of misspecification, both MLE and the robust estimator yield accurate estimates. Both estimates are nearly equivalent to one another in the sense that their point estimates, standard deviation, and coverage at significance level $\alpha=0.05$ are very similar (Table 1). However, when misspecification is introduced, MLE, sample correlation, and robust estimator yield noticeably different results. At misspecification degree $\varepsilon=0.1$, MLE and sample correlation are substantially biased with average estimates of $0.097$ and $0.084$, corresponding to biases of $-0.403$ and $-0.416$, respectively, as well as zero coverage. In contrast, the robust estimator maintains accuracy with an average estimate of $0.466$, which corresponds to only a minor bias of $-0.034$ and a good coverage of $0.911$ (Table 1). When the misspecification is increased to $\varepsilon=0.2$, the contrast between the estimators becomes even stronger. While the robust estimator is still remarkably close to the truth with a small bias of $-0.061$, MLE and sample correlation produce estimates that are not only severely biased (biases of $-0.672$ and $-0.633$), but also sign-flipped: While the true correlation is strongly positive (0.5), both estimates are considerably negative ($-0.172$ and $-0.133$). It is worth noting that in the presence of misspecification, the confidence intervals of the robust estimator are wider than those of the MLE (see Table 1). This is expected because of a well-known trade-off between robustness and efficiency: An estimator that is designed to reduce bias, like a robust estimator, will inevitably have a larger estimation variance (e.g., Huber & Ronchetti,, 2009). These wider confidence intervals furthermore explain why the robust estimator improves its coverage in Table 1 when the misspecification fraction is increased from 0.1 to 0.2.
6.2 Polychoric correlation matrix
The goal of this simulation study is to robustly estimate a polychoric correlation matrix, that is, a matrix comprising of pairwise polychoric correlation coefficients. The simulation design is based on Foldnes & Grønneberg, (2020).
Let there be $q$ observed ordinal random variables where a latent variable underlies each ordinal variable. In accordance with the multivariate polychoric model (e.g., Muthén,, 1984), the latent variables are assumed to jointly follow a $q$-dimensional normal distribution with mean zero and a covariance matrix with unit diagonal elements so that the covariance matrix is a correlation matrix. Each individual latent variable is discretized to its corresponding observed ordinal variable akin to discretization process (1).
Variable 1 2 3 4 5
1 1.00
2 0.56 1.00
3 0.48 0.42 1.00
4 0.40 0.35 0.30 1.00
5 0.32 0.28 0.24 0.20 1.00
Table 2: Correlation matrix of $q=5$ latent variables as in Foldnes & Grønneberg, (2020). In line with the multivariate polychoric correlation model (e.g., Muthén,, 1984), the latent variables are jointly normally distributed with mean zero and this correlation matrix as covariance matrix.
Following the five-dimensional simulation design in Foldnes & Grønneberg, (2020), there are $q=5$ ordinal variables with polychoric correlation matrix as in Table 5 such that the pairwise correlations vary from a low 0.2 to a moderate 0.56.333The correlation matrix in Table 2 has the additional interpretation of being the covariance matrix of a factor model for a single factor with loadings vector $(0.8,0.7,0.6,0.5,0.4)^{\top}$. For all latent variables, the discretization thresholds are set to, in ascending order, $\Phi_{1}^{-1}(0.1)=-1.28,\ \Phi_{1}^{-1}(0.3)=-0.52,\ \Phi_{1}^{-1}(0.7)=0.56$, and $\Phi_{1}^{-1}(0.9)=1.28$, such that each ordinal variable can take five possible values. A visualization of the implied distribution of each ordinal variable can be found in Figure 5 in Foldnes & Grønneberg, (2020).
As misspecifying distribution, we choose a $q$-dimensional Gumbel distribution comprising of mutually independent Gumbel marginal distributions, each with location and scale parameters equal to 0 and 3, respectively. To obtain ordinal observations, the unobserved realizations from this distribution are discretized via the same threshold values as realizations from the model (normal) distribution. As such, the uninformative ordinal observations generated by this misspecifying distribution emulate the erratic behavior of a careless respondent: Unlike in the previous simulation design (Section 6.1), the uninformative responses are not concentrated around a few response options, but may occur in every response option.
For misspecification fraction $\varepsilon\in\{0,0.1,0.2\}$, we sample $N=1,000$ ordinal five-dimensional responses from this data generating process and use them to estimate the polychoric correlation matrix in Table 2 via ML and our robust estimator (again with tuning constant $c=1.6$). We repeat this procedure 1,000 times.
Figure 5: Absolute average bias of the ML estimator (top) and our robust estimator (bottom) for each unique pairwise polychoric correlation coefficient in the true correlation matrix (Table 2), expressed as a function of the misspecification fraction $\varepsilon$ ($x$-axis). Averages are taken over 1,000 repetitions of the data generating process described in Section 6.2.
Figure 5 visualizes the absolute average bias (calculated over the 1,000 repetitions) of the ML and robust estimator for each pairwise polychoric correlation coefficient in Table 2. As expected, when the model is correctly specified ($\varepsilon=0$), then both estimators coincide with accurate estimates. However, in the presence of partial misspecification ($\varepsilon>0$), the two estimators deviate. The ML estimator exhibits a notable bias for all correlation coefficient, which increases with increasing misspecification fraction. The magnitude of the bias tends to be larger for pairs with larger true correlation, such as $0.56$ for pair $(1,2)$, than for pairs with weaker true correlation, such as $0.20$ for pair $(5,4)$. Conversely, the robust estimator remains nearly unaffected throughout all considered misspecification fractions for each correlation coefficient, reflecting an excellent performance with respect to robustness to uninformative responses.
The two simulation studies demonstrate that already a small degree of partial misspecificaion to uninformative responses, such as careless responses, can render the commonly employed ML estimator unreliable, while the proposed robust estimator retains good accuracy even in the presence of a considerable number of uninformative responses. On the other hand, when the model is correctly specified, the ML estimator and robust estimator produce equivalent estimates.
7 Empirical application
7.1 Background and study design
We demonstrate our proposed method on empirical data by using a subset of the 100 unipolar markers of the Big Five personality traits (Goldberg,, 1992). Each marker is a an item comprising a single English adjective (such as “bold” or “timid”) asking respondents to indicate how accurately the adjective describes their personality using a 5-point Likert-type rating scale (very inaccurate, moderately inaccurate, neither accurate nor inaccurate, moderately accurate, and very accurate). Here, each Big Five personality trait is measured with six pairs of adjectives that are polar opposites to one another (such as “talkative” vs. “silent”), that is, twelve items in total for each trait. It seems implausible that an attentive respondent would choose to agree (or disagree) to both items in a pair of polar opposite adjectives. Consequently, one would expect a strongly negative correlation between polar adjectives if all respondents respond attentively (Arias et al.,, 2020).
Arias et al., (2020) collect measurements of three Big Five traits in this way, namely extroversion, neuroticism, and conscientiousness. [^4] Arias et al., (2020) synonymously refer to neuroticism as emotional stability. Furthermore, in addition to the three listed traits, Arias et al., (2020) collect measurements of the trait dispositional optimism by using a different instrument, and another scale that is designed to not measure any construct. We do not consider these scales in this empirical demonstration. The sample that we shall use, Sample 1 in Arias et al., (2020), consists of $N=725$ online respondents who are all U.S. citizens, native English speakers, and tend to have relatively high levels of reported education (about 90% report to hold an undergraduate or higher degree). Concerned about respondent inattention in their data, Arias et al., (2020) construct a factor mixture model for detecting inattentive/careless participants. Their model crucially relies on response inconsistencies to polar opposite adjectives and is designed to primarily detect careless straightlining responding. They find that careless responding is a sizable problem in their data. Their model finds evidence of straightliners, and the authors conclude that if unaccounted for, they can substantially deteriorate the fit of theoretical models, produce spurious variance, and overall jeopardize the validity of research results.
Due to the suspected presence of careless respondents, we apply our proposed method to estimate the polychoric correlation coefficients between all $\binom{12}{2}=66$ unique item pairs in the neuroticism scale to obtain an estimate of the scale’s (polychoric) correlation matrix. The results of the remaining two scales are qualitatively similar and are reported in Appendix C. We estimate the polychoric correlation matrix twice: via classic maximum likelihood and via our proposed robust alternative, with tuning parameter $c=1.6$. The results remain qualitatively similar for different finite choices of $c$.
7.2 Results
Figure 6: Difference between absolute estimates for the polychoric correlation coefficient of our robust estimator and the MLE for each item pair in the neuroticism scale, using the data of Arias et al., (2020). The items are “calm” (N1_P), “angry” (N1_N), “relaxed” (N2_P), “tense” (N2_N), “at ease” (N3_P), “nervous” (N3_N), “not envious” (N4_P), “envious” (N4_N), “stable” (N5_P), “unstable” (N5_N), “contented” (N6_P), and “discontented” (N6_N). For the item naming given in parentheses, items with identical identifier (the integer after the first “N”) are polar opposites, where a last character “P” refers to the positive opposite and “N” to the negative opposite. The individual estimates of each method are provided in Table C.2 in Appendix C.
Sample cor MLE Robust
Parameter Estimate SE Estimate SE Estimate SE
$\rho$ $-0.562$ 0.031 $-0.618$ 0.025 $-0.925$ 0.062
$a_{1}$ $-1.370$ 0.061 $-1.570$ 0.276
$a_{2}$ $-0.476$ 0.043 $-0.560$ 0.203
$a_{3}$ $0.121$ 0.042 $0.109$ 0.187
$a_{4}$ $1.060$ 0.054 $1.080$ 0.105
$b_{1}$ $-0.857$ 0.049 $-0.905$ 0.073
$b_{2}$ $-0.004$ 0.041 $-0.040$ 0.091
$b_{3}$ $0.608$ 0.045 $0.640$ 0.364
$b_{4}$ $1.580$ 0.071 $1.171$ 0.811
Table 3: Parameter estimates with standard errors (SEs) for the correlation between the neuroticism adjective pair “envious” and “not envious” in the data of Arias et al., (2020), using Pearson sample correlation, ML, and robust estimator with tuning constant $c=1.6$. Each adjective item has five Likert-type answer categories. The sample correlation coefficient does not model thresholds, hence no estimates can be reported for them.
Figure 6 visualizes the difference in absolute estimates for the polychoric correlation coefficient between all 66 unique item pairs in the neuroticism scale. For all unique pairs, our method estimates a stronger correlation coefficient than ML. The differences in absolute estimates on average amount to 0.083, ranging from only marginally larger than zero to a substantive 0.314. For correlations between polar opposite adjectives, the average absolute difference between our robust method and ML is 0.151. The fact that a robust method consistently yields stronger correlation estimates than ML, particularly between polar opposite adjectives, is indicative of the presence of negative leverage points, which drag negative correlation estimates towards zero, that is, they attenuate the estimated strength of correlation. Here, such negative leverage points could be the responses of careless respondents who report agreement or disagreement to both items in item pairs that are designed to be negatively correlated. For instance, recall that it is implausible that an attentive respondent would choose to agree (or disagree) to both adjectives in the pair “envious” and “not envious” (cf., Arias et al.,, 2020). If sufficiently many such respondents are present, then the presumably strongly negative correlation between these two opposite adjectives will be estimated to be weaker than it actually is.
To further investigate the presence of careless respondents who attenuate correlational estimates, we study in detail the adjective pair “not envious” and “envious”, which featured the largest discrepancy between the ML estimate and robust estimate in Figure 6, with an absolute difference of 0.314. The results of the two estimators and, for completeness, the sample correlation, are summarized in Table 3. The ML estimate of $-0.618$ and sample correlation estimate of $-0.562$ for the (polychoric) correlation coefficient seem remarkably weak considering that the two adjectives in question are polar opposites. In contrast, the robust correlation estimate is given by $-0.925$, which seems much more in line with what one would expect if all participants responded accurately and attentively (cf., Arias et al.,, 2020).
$X$\$Y$ 1 2 3 4 5
1 0.019 0.007 0.003 0.028 0.022
2 0.007 0.040 0.050 0.138 0.014
3 0.006 0.047 0.143 0.030 0.003
4 0.054 0.189 0.029 0.019 0.007
5 0.108 0.018 0.006 0.008 0.007
(a) Empirical relative frequencies $\widehat{f}_{N}(x,y)$
$X$\$Y$ 1 2 3 4 5
1 $<0.001$ $<0.001$ $<0.001$ 0.024 0.034
2 $<0.001$ 0.004 0.062 0.153 0.010
3 0.001 0.072 0.145 0.038 $<0.001$
4 0.061 0.205 0.047 0.002 $<0.001$
5 0.120 0.020 $<0.001$ $<0.001$ $<0.001$
(b) Estimated response probabilities $p_{xy}(\widehat{\bm{\theta}}_{N})$
$X$\$Y$ 1 2 3 4 5
1 $>1,000$ $>1,000$ 11.82 1.14 0.65
2 $>1,000$ 10.07 0.80 0.90 1.42
3 5.48 0.65 0.99 0.80 77.14
4 0.88 0.92 0.61 12.66 $>1,000$
5 0.89 0.88 36.01 $>1,000$ $>1,000$
(c) DPRs $\widehat{f}_{N}(x,y)\big{/}p_{xy}(\widehat{\bm{\theta}}_{N})$
Table 4: Empirical relative frequency (top), estimated response probability (center), and decentralized Pearson residual (bottom) of each response $(x,y)$ for the “not envious” ($X$)–“envious” ($Y$) item pair in the measurements of Arias et al., (2020) of the neuroticism scale. Estimate $\widehat{\bm{\theta}}_{N}$ was computed with tuning constant $c=1.6$. The complete DPR values are provided in Table C.5 in the appendix.
To study the potential presence of careless responses in each contingency table cell $(x,y)$ for item pair “envious” and “not envious”, Table 4 lists the DPRs at the robust estimate, alongside the associated model probabilities and empirical relative frequencies. [^5] A visualization of Table 4 is provided in Figure C.5 in the appendix. A total of six cells have extremely large DPR values of higher than 1,000, and, in addition, five cells have a DPR of higher than 10, and one cell has a DPR of higher than 5. Such DPR values are far away from ideal value 1 at which the model would fit perfectly, thereby indicating a poor fit of the polychoric model for such responses. It stands out that all such poorly fitted cells are those whose responses might be viewed as inconsistent. Indeed, response cells $(x,y)=(1,1),(1,2),(2,1),(2,2)$ indicate that a participant reports that neither “envious” nor “not envious” characterizes them accurately, which are mutually contradicting responses, while for response cells $(x,y)=(4,4),(4,5),(5,4),(5,5)$ both adjectives characterize them accurately, which is again contradicting. As discussed previously, such responses might due to careless responding. The robust estimator suggests that such responses cannot be fitted well by the polychoric model and subsequently downweighs their influence in the estimation procedure by mapping their Pearson residual with the linear part of the discrepancy function $\varphi(\cdot)$ in (11). Notably, also cells $(x,y)=(1,3),(3,1),(3,5),(5,3)$ are poorly fitted. These responses report (dis)agreement to one opposite adjective, while being neutral about the other opposite. It is beyond the scope of this paper to assess whether such response patterns are also indicative of careless responding, but the robust estimator suggests that such responses at least cannot be fitted well by the polychoric model with the data of Arias et al., (2020).
Overall, leveraging our robust estimator, we find evidence for the presence of careless respondents in the data of Arias et al., (2020). While they substantially affect the correlation estimate of ML, amounting to about $-0.62$, which is much weaker than one would expect for polar opposite items, our robust estimator can withstand their influence with an estimate of about $-0.93$ and also help identify them through unreasonably large DPR values.
8 Estimation under distributional misspecification
8.1 Preliminaries
The simulation studies in Section 6 have been concerned with a situation in which the polychoric model is misspecified for a subset of a sample (partial misspecification). However, as outlined in Section 4.3, another misspecification framework of interest is that of distributional misspecification where the model is misspecified for the entire sample. Suppose that instead of a bivariate standard normal distribution, the latent variables $(\xi,\eta)$ jointly follow an unknown and unspecified distribution $G$. In this framework, the object of interest is the population correlation between $\xi$ and $\eta$ under distribution $G$, that is, $\rho_{G}=\mathrm{Cor}_{G}\left[\xi,\ \eta\right]$, rather than a polychoric correlation coefficient.
Although distributional misspecification is not covered by the misspecified distribution (7) under which we study the theoretical properties of our proposed estimator, our estimator could still offer a gain in robustness compared to ML estimation under distributional misspecification. Specifically, the robust estimator may still be useful if the central part of the nonnormal distribution $G$ is not too different from a standard bivariate normal distribution: Intuitively, if the difference between $G$ and a standard normal distribution is mainly in the tails, $G$ can be approximated by a mixture of a standard normal distribution (to cover the central part) and some other distribution $H$ (to cover the tails). The ML estimator tries to treat influential observations from the tails—which cannot be fitted well by the polychoric model—as if they were normally distributed, resulting in a possibly large estimation bias. In contrast, the robust estimator uses the normal distribution only for observations from the central part—which may fit the polychoric model well enough—and downweights observations from the tails. Thus, as long as such a mixture is a decent approximation of the nonnormal distribution $G$, the robust estimator should perform reasonably well. However, if $G$ cannot be approximated by such a mixture, neither the ML estimator nor our estimator can be expected to perform well. Overall, though, our estimator could offer an improvement in terms of robustness to distributional misspecification.
In the following, we perform a simulation study to investigate the performance of our estimator when the polychoric model is distributionally misspecified.
8.2 Simulation study
To simulate ordinal variables that were generated by a nonnormal latent distribution $G$, we employ the VITA simulation method of Grønneberg & Foldnes, (2017). For a pre-specified value of the population correlation $\rho_{G}$, the VITA method models the latent random vector $(\xi,\eta)$ such that the individual variables $\xi$ and $\eta$ both possess standard normal marginal distributions with population correlation set equal to $\rho_{G}=\mathrm{Cor}_{G}\left[\xi,\ \eta\right]$, but are not jointly normally distributed. Instead, their joint distribution $G$ is equal to a pre-specified nonnormal copula distribution, such as the Clayton or Gumbel copula. Grønneberg & Foldnes, (2017) show that discretizing such VITA latent variables yields ordinal observations that could have not been generated by a standard bivariate normal distribution, thereby ensuring proper violation of the latent normality assumption.
Figure 7: Bivariate density plots of the standard normal distribution (left) and Clayton copula with standard normal marginals (right), for population correlations 0.9 (top) and 0.3 (bottom).
Figure 8: Boxplot visualization of the bias of the robust and ML estimator for the polychoric correlation coefficient, $\widehat{\rho}_{N}-\rho_{G}$, under distributional misspecification through a Clayton copula with correlation $\rho_{G}=0.9$ (left) and $\rho_{G}=0.3$ (right), across 1,000 repetitions. The tuning constant of the robust estimator is set to $c=1.6$.
To investigate the robustness of our estimator to distributional misspecification, we use the VITA method to generate draws of the latent variables $(\xi,\eta)$ such that the latent variables are jointly distributed according to a Clayton copula $G$ with population correlation $\rho_{G}\in\{0.9,0.3\}$ (see Figure 7 for visualizations). Following the discretization process (1), we discretize both latent variables via discretization thresholds
such that both resulting ordinal variables have five response options each. We generate $N=1,000$ ordinal responses according to this data generation process and compute across 1,000 repetitions the same estimators and performance measures as in the simulations in Section 6.
Figure 8 provides a boxplot visualization of the bias of each estimator under both Clayton copulas across the repetitions. [^6] In seven of the 1,000 repetitions for the Clayton copula with correlation 0.9, the robust estimator experienced numerical instability and subsequently did not converge to a solution. Consequently, these unconverged estimates are omitted from Figure 8. We explain this numerical issue in detail in Appendix B. For correlation 0.9, the ML estimator exhibits a noteworthy bias, whereas the robust estimator remains accurate, albeit with a larger estimation variance as compared to partial misspecification (cf., Section 6). Conversely, for the weaker correlation 0.3, both estimators are fairly accurate with median biases of about 0.025.
These results suggest that the Clayton copula with correlation 0.9 might be reasonably approximable by a mixture of a normal distribution that covers the center of the probability mass and some nonnormal distribution that covers the tails (cf., Section 8.1). Indeed, Figure 7 indicates that the normal distribution and Clayton copula at correlation 0.9 seem to behave similarly in the center, but deviate from one another towards the tails. On the other hand, at correlation 0.3, the two densities do not appear to be drastically different from one another, which may explain why both ML and robust estimator work reasonably well for such distributional misspecification: If the latent distribution is not too different from a normal distribution, then the polychoric model may offer a satisfactory fit despite being technically misspecified.
Overall, this simulation study demonstrates that in some cases of distributional misspecification, robustness can be gained with our estimator, compared to ML estimation. However, it also demonstrates that there are cases of distributional misspecification for which the ML estimator still works quite well such that the robust estimator offers little gain. Nevertheless, the fact that in some situations robustness can be gained under distributional misspecification represents an overall gain in robustness.
9 Discussion and conclusion
We consider a situation where the polychoric correlation model is potentially misspecified for a subset of responses in a sample, that is, a set of uninformative observations not generated by a latent standard normal distribution. This model misspecification framework, called partial misspecification here, stems from the robust statistics literature, and a relevant special case is that of careless respondents in questionnaire studies. We demonstrate that maximum likelihood estimation is highly susceptible to the presence of such uninformative responses, resulting in possibly large estimation biases. As a remedy, we propose an estimator based on Welz, (2024) that is designed to be robust to partial model misspecification. Our estimator generalizes maximum likelihood estimation, does not make any assumption on the magnitude or type of potential misspecification, comes at no additional computational cost, and possesses attractive statistical guarantees such as asymptotic normality. It furthermore allows to pinpoint the sources of potential model misspecification through the notion of decentralized Pearson residuals (DPRs). Each possible response option is assigned a DPR, where values substantially larger than the ideal value 1 imply that the response in question cannot be fitted well by the polychoric correlation model. In addition, the methodology proposed in this paper is implemented in the free open source package robcat for the statistical programming environment R and is publicly available at https://github.com/mwelz/robcat.
Although not covered by our partial misspecification framework, we also discuss how and when our estimator can offer a robustness gain (compared to ordinary maximum likelihood) when the polychoric model is misspecified for all observations in a sample, which has been a subject of interest in recent literature. In essence, there can be a robustness gain if the latent nonnormal distribution that generated the data can be reasonably well approximated by a mixture of a normal distribution that covers the center of probability mass and some unspecified nonnormal distribution that covers the tails.
We verify the enhanced robustness and theoretical properties of our robust generalized maximum likelihood estimator in simulation studies. Furthermore, we demonstrate the estimator’s practical usefulness in an empirical application on a Big Five administration, where we find compelling evidence for the presence of careless respondents as source of partial model misspecification.
However, our estimator depends on a user-specified choice of a tuning constant $c$, which governs a tradeoff between robustness and efficiency. While simulation experiments suggest that the choice $c=1.6$ provides a good tradeoff and estimates do not change considerably for other finite choices of $c$, a detailed investigation on this tuning constant needs to be carried out in future work.
The methodology proposed in this paper suggests a number of extensions. For instance, one could use a robustly estimated polychoric correlation matrix in structural equation models to robustify such models against misspecification. Similar robustification could by achieved in, for instance but not limited to, principal component analyses, multidimensional scaling, or clustering. In addition, Welz, (2024) proposes a general statistical test to pinpoint possible sources of model misspecification on the individual response level. This test could be used to test in a statistically sound way whether a given response can be fitted well by the polychoric correlation model. We leave these avenues to further research.
Overall, our novel robust estimator could open the door for a new line of research that is concerned with making the correlation-based analysis of rating data more reliable by reducing dependence on modeling assumptions.
Appendix A Asymptotic properties of the robust estimator
This appendix section presents the limit theory of the proposed robust estimator $\widehat{\bm{\theta}}_{N}$ in (10). Throughout this section, we assume that the number of response categories, $K_{X}$ and $K_{Y}$, are fixed and known, and that the sample $\{(X_{i},Y_{i})\}_{i=1}^{N}$ used to compute an estimate $\widehat{\bm{\theta}}_{N}$ has been generated by the process in (1) where the latent $\{(\xi_{i},\eta_{i})\}_{i=1}^{N}$ are draws from distribution $G_{\varepsilon}$ in (7) with unobserved misspecification fraction $\varepsilon\in[0,0.5)$ and unknown misspecifying distribution $H$.
The following three subsections first state the main theorem, then discuss its assumptions, and then provide closed-form expressions of quantities used in the asymptotic analysis.
A.1 Main theorem
We start by introducing additional notation.
For fixed tuning constant $c\geq 1$, let
$w(z)=\mathds{1}\left\{z\in[0,c]\right\}+c\mathds{1}\left\{z>c\right\}/z\quad\mathrm{for}\ z\geq 0,$
with first derivative
$w^{\prime}(z)=0\mathds{1}\left\{z\in[0,c]\right\}-c\mathds{1}\left\{z>c\right\}/z^{2}.$
Define the $d$-dimensional gradient of $\log\left(p_{xy}(\bm{\theta})\right)$ for response $(x,y)\in\mathcal{X}\times\mathcal{Y}$ at parameter $\bm{\theta}\in\bm{\Theta}$ as
$\bm{s}_{xy}(\bm{\theta})=\frac{1}{p_{xy}(\bm{\theta})}\left(\frac{\partial}{\partial\bm{\theta}}\ p_{xy}(\bm{\theta})\right),$
where one can write for the gradient of $p_{xy}(\bm{\theta})$
$\frac{\partial}{\partial\bm{\theta}}\ p_{xy}(\bm{\theta})=\frac{\partial}{\partial\bm{\theta}}\Phi_{2}\left(a_{x},b_{y};\rho\right)-\frac{\partial}{\partial\bm{\theta}}\Phi_{2}\left(a_{x-1},b_{y};\rho\right)-\frac{\partial}{\partial\bm{\theta}}\Phi_{2}\left(a_{x},b_{y-1};\rho\right)+\frac{\partial}{\partial\bm{\theta}}\Phi_{2}\left(a_{x-1},b_{y-1};\rho\right),$
see, e.g., Olsson, (1979, equation 4), and define the $d\times d$ Hessian matrix of $\log\left(p_{xy}(\bm{\theta})\right)$ as
$\bm{Q}_{xy}(\bm{\theta})=\frac{1}{p_{xy}(\bm{\theta})}\left(\frac{\partial^{2}}{\partial{\bm{\theta}}\partial{\bm{\theta}}^{\top}}p_{xy}(\bm{\theta})\right)-\bm{s}_{xy}(\bm{\theta})\bm{s}_{xy}(\bm{\theta})^{\top}.$ (A.1)
We derive closed-form expressions of all components in $\bm{Q}_{xy}(\bm{\theta})$ in Subsection A.3. In addition, for $K=K_{X}\cdot K_{Y}$, the total number of contingency table cells and $d$-dimensional vectors
$\bm{w}_{xy}\left(\bm{\theta}\right)=\bm{s}_{xy}(\bm{\theta})\mathds{1}\left\{\frac{f_{\varepsilon}\left(x,y\right)}{p_{xy}(\bm{\theta})}\in[0,c]\right\},$
define the $d\times K$ matrix
$\bm{W}(\bm{\theta})=$
$\quad\bigg{(}\bm{w}_{11}\left(\bm{\theta}\right),\cdots,\bm{w}_{1,K_{Y}}\left(\bm{\theta}\right),\bm{w}_{21}\left(\bm{\theta}\right),\cdots,\bm{w}_{2,K_{Y}}\left(\bm{\theta}\right),\cdots,\bm{w}_{K_{X},1}\left(\bm{\theta}\right),\bm{w}_{K_{X},2}\left(\bm{\theta}\right),\cdots,\bm{w}_{K_{X},K_{Y}}\left(\bm{\theta}\right)\bigg{)}$
that row-binds all $K$ vectors $\bm{s}_{xy}(\bm{\theta})$ multiplied by an indicator that takes value 1 when associated population DPR is in the MLE-part of the function $\varphi(\cdot)$ in (11) and 0 otherwise. In similar fashion, define the $K$-dimensional vector
$\bm{f}_{\varepsilon}=\Big{(}f_{\varepsilon}\left(1,1\right),\dots,f_{\varepsilon}\left(1,K_{Y}\right),f_{\varepsilon}\left(2,1\right),\dots,f_{\varepsilon}\left(2,K_{Y}\right),\dots,f_{\varepsilon}\left(K_{X},1\right),f_{\varepsilon}\left(K_{X},2\right),\dots,f_{\varepsilon}\left(K_{X},K_{Y}\right)\Big{)}^{\top}$
that holds all $K$ evaluations of the function $f_{\varepsilon}$, and put
$\bm{\Omega}=\mathrm{diag}(\bm{f}_{\varepsilon})-\bm{f}_{\varepsilon}\bm{f}_{\varepsilon}^{\top},$
where $\mathrm{diag}(\bm{f}_{\varepsilon})$ is a $K\times K$ diagonal matrix that holds the coordinates of $\bm{f}_{\varepsilon}$ on its principal diagonal.
The following theorem establishes consistency and asymptotic normality of the estimator. This theorem follows immediately from Theorems 1 and 2 in Welz, (2024).
Theorem 1.
Under certain regularity conditions that do not restrict the degree or type of possible misspecification of the polychoric model beyond $\varepsilon\in[0,0.5)$, when $N\to\infty$ it holds true that
$\widehat{\bm{\theta}}_{N}\stackrel{{\scriptstyle\mathbb{P}}}{{\longrightarrow}}\bm{\theta}_{0},$
as well as
$\sqrt{N}\left(\widehat{\bm{\theta}}_{N}-\bm{\theta}_{0}\right)\stackrel{{\scriptstyle\mathrm{d}}}{{\longrightarrow}}\textnormal{N}_{d}\Big{(}\bm{0},\bm{\Sigma}\left({\bm{\theta}_{0}}\right)\Big{)},$
where, as a function of $\bm{\theta}\in\bm{\Theta}$, the estimator’s invertible asymptotic covariance matrix is given by
$\bm{\Sigma}\left({\bm{\theta}}\right)=\bm{M}\left(\bm{\theta}\right)^{-1}\bm{U}\left(\bm{\theta}\right)\bm{M}\left(\bm{\theta}\right)^{-1},$
where
$\bm{U}\left(\bm{\theta}\right)$ $=\bm{W}\left(\bm{\theta}\right)\bm{\Omega}\bm{W}\left(\bm{\theta}\right)^{\top}\qquad\text{and}$
$\bm{M}\left(\bm{\theta}\right)$ $=\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}f_{\varepsilon}\left(x,y\right)\Bigg{(}{w^{\prime}}\Bigg{(}\frac{f_{\varepsilon}\left(x,y\right)}{p_{xy}(\bm{\theta})}\Bigg{)}\frac{f_{\varepsilon}\left(x,y\right)}{p_{xy}(\bm{\theta})}\bm{s}_{xy}(\bm{\theta})\bm{s}_{xy}(\bm{\theta})^{\top}-{w}\Bigg{(}\frac{f_{\varepsilon}\left(x,y\right)}{p_{xy}(\bm{\theta})}\Bigg{)}\bm{Q}_{xy}(\bm{\theta})\Bigg{)}$
are $d\times d$ symmetric and invertible matrices.
We list and discuss the regularity conditions in Subsection A.2.
It can be shown that in the absence of misspecification (such that $\bm{\theta}_{0}=\bm{\theta}_{*}$), the asymptotic covariance matrix $\bm{\Sigma}\left({\bm{\theta}_{*}}\right)$ is equal to the inverted Fisher information matrix of the polychoric model (Lemma B.5 in Welz,, 2024), which is well-known to be the asymptotic covariance matrix of the MLE. It follows that under correct specification of the polychoric model, the generalized ML estimator is indeed asymptotically first and second order equivalent to the consistent and efficient MLE of Olsson, (1979).
A consistent estimator of the unobserved asymptotic covariance matrix $\bm{\Sigma}\left({\bm{\theta}_{0}}\right)$ can be constructed as follows. Replace all population class probabilities $f_{\varepsilon}\left(x,y\right)$ by their corresponding empirical counterparts $\widehat{f}_{N}(x,y)$ in matrices $\bm{W}\left(\bm{\theta}\right),\bm{M}\left(\bm{\theta}\right)$, and $\bm{\Omega}$. Then exploit the plug-in principle and evaluate $\bm{U}\left(\bm{\theta}\right)$ and $\bm{M}\left(\bm{\theta}\right)$ at the point estimate $\widehat{\bm{\theta}}_{N}.$ Denote the ensuing plug-in estimator by $\bm{\Sigma}\left({\widehat{\bm{\theta}}_{N}}\right)$, which is consistent for $\bm{\Sigma}\left({\bm{\theta}_{0}}\right)$ by Theorem 1 and the continuous mapping theorem.
A.2 Assumptions
In the following, we list and discuss a set of assumptions that constitute the regularity conditions in Theorem 1. These assumptions are based on Welz, (2024).
Assumption Set A.
Suppose that the following assumptions hold true.
  1. A.1
    $c\in[1,+\infty]$,
  2. A.2
    $\bm{\Theta}\subset\mathbb{R}^{d}$ is compact, where $d=K_{X}+K_{Y}-1$ denotes the number of parameters in the polychoric model,
  3. A.3
    $\bm{\theta}_{0}=\arg\min_{\bm{\theta}\in\bm{\Theta}}L(\bm{\theta},f_{\varepsilon})$ is a unique minimum,
  4. A.4
    #$\#\big{\{}(x,y)\in\mathcal{X}\times\mathcal{Y}:f_{\varepsilon}\left(x,y\right)>0\big{\}}>d$,
  5. A.5
    $\frac{f_{\varepsilon}\left(x,y\right)}{p_{xy}(\bm{\theta}_{0})}\neq c$ for all $(x,y)\in\mathcal{X}\times\mathcal{Y}$.
Assumptions A.1A.3 are required for consistency of $\widehat{\bm{\theta}}_{N}$, whereas all assumptions are required for its asymptotic normality in Theorem 1. Under these assumptions, the theorem follows immediately by combining Theorems 1 and 2 in Welz, (2024).iiiThe general framework of Welz, (2024) entertains a total of nine assumptions instead of the five assumptions listed here. The four remaining assumptions are automatically satisfied in the polychoric model and are therefore omitted.
Assumption A.1 ensures that the discrepancy function $\varphi(\cdot)$ exhibits meaningful behavior when evaluated at DPRs, such as the ideal residual value 1 being included in the interval $[0,c]$. Compactness of the parameter space (Assumption A.2) is primitive and required for a technicality when proving consistency of $\widehat{\bm{\theta}}_{N}$ (see the proof of Theorem 1 in Welz,, 2024, for details). Uniqueness of a global minimum (Assumption A.3) is a common assumption in the literature on $M$-type estimators (e.g., Chapter 5.2 in Van der Vaart,, 1998). Assumption A.4 imposes that the number of positive response population probabilities is strictly larger than the number of model parameters. In other words, there must be more sources of variation (populated contingency table cells) than parameters. One may view this assumption as a rank condition that ensures invertibility of the Hessian matrix of the minimization problem (9), and is required to prevent rank deficiency of the asymptotic covariance matrix of estimator $\widehat{\bm{\theta}}_{N}$. Since $\widehat{f}_{N}(x,y)>0$ implies $f_{\varepsilon}\left(x,y\right)>0$, this assumption can be empirically verified. Finally, Assumption A.5 requires that no DPR at the global minimum is equal to tuning constant $c$. This is a primitive condition that is required for the loss function to be twice differentiable at $\bm{\theta}_{0}$, which is a requirement for the existence of the estimator’s asymptotic covariance matrix.
Finally, observe that no assumption is made that restricts the degree or type of potential misspecification of the polychoric model.
A.3 Expressions of first and second order derivatives
This section presents closed-form expressions of all components of matrix $\bm{Q}_{xy}(\bm{\theta})$ in (A.1).
A.3.1 First order terms
For response $(x,y)\in\mathcal{X}\times\mathcal{Y}$ and $\bm{\theta}\in\bm{\Theta}$, the gradient of $p_{xy}(\bm{\theta})$ can be expressed as
$\frac{\partial p_{xy}(\bm{\theta})}{\partial\bm{\theta}}=\frac{\partial}{\partial\bm{\theta}}\Phi_{2}\left(a_{x},b_{y};\rho\right)-\frac{\partial}{\partial\bm{\theta}}\Phi_{2}\left(a_{x-1},b_{y};\rho\right)-\frac{\partial}{\partial\bm{\theta}}\Phi_{2}\left(a_{x},b_{y-1};\rho\right)+\frac{\partial}{\partial\bm{\theta}}\Phi_{2}\left(a_{x-1},b_{y-1};\rho\right),$ (A.2)
see, e.g., Olsson, (1979, equation 4). To characterize this gradient, we provide expressions for individual partial derivatives of $p_{xy}(\bm{\theta})$, that is,
$\frac{\partial p_{xy}(\bm{\theta})}{\partial\bm{\theta}}=\left(\frac{\partial p_{xy}(\bm{\theta})}{\partial\rho},\frac{\partial p_{xy}(\bm{\theta})}{\partial a_{1}},\dots,\frac{\partial p_{xy}(\bm{\theta})}{\partial a_{K_{X}-1}},\frac{\partial p_{xy}(\bm{\theta})}{\partial b_{1}},\dots,\frac{\partial p_{xy}(\bm{\theta})}{\partial b_{K_{Y}-1}}\right)^{\top}.$
First, for any $u,v\in\mathbb{R}$, it can be shown (e.g., Drezner & Wesolowsky,, 1990) that
$\frac{\partial}{\partial\rho}\Phi_{2}\left(u,v;\rho\right)=\phi_{2}\left(u,v;\rho\right),$
as well as (e.g., Tallis,, 1962)
$\frac{\partial}{\partial u}\Phi_{2}\left(u,v;\rho\right)=\phi_{1}\left(u\right)\Phi_{1}\left(\frac{v-\rho u}{\sqrt{1-\rho^{2}}}\right),$
where $\phi_{1}\left(\cdot\right)$ and $\Phi_{1}\left(\cdot\right)$ denote the density and distribution function, respectively, of the univariate standard normal distribution. The complementary partial derivative with respect to $v$ follows analogously by symmetry.
It now follows immediately from (A.2) that the partial derivative of $p_{xy}(\bm{\theta})$ with respect to $\rho$ is given by
$\frac{\partial p_{xy}(\bm{\theta})}{\partial\rho}=\phi_{2}\left(a_{x},b_{y};\rho\right)-\phi_{2}\left(a_{x-1},b_{y};\rho\right)-\phi_{2}\left(a_{x},b_{y-1};\rho\right)+\phi_{2}\left(a_{x-1},b_{y-1};\rho\right),$
whereas the partial derivatives with respect to the individual thresholds are characterized by
$\frac{\partial p_{xy}(\bm{\theta})}{\partial a_{k}}=\begin{dcases}\frac{\partial}{\partial a_{x}}\Phi_{2}\left(a_{x},b_{y};\rho\right)-\frac{\partial}{\partial a_{x}}\Phi_{2}\left(a_{x},b_{y-1};\rho\right)&\text{if }k=x,\\ -\frac{\partial}{\partial a_{x-1}}\Phi_{2}\left(a_{x-1},b_{y};\rho\right)+\frac{\partial}{\partial a_{x-1}}\Phi_{2}\left(a_{x-1},b_{y-1};\rho\right)&\text{if }k=x-1,\\ 0&\text{otherwise},\end{dcases}$
for $k=1,\dots,K_{X}-1$. An expression for $\frac{\partial p_{xy}(\bm{\theta})}{\partial b_{k}}$ can be derived analogously.
A.3.2 Second order terms
Here we provide expressions for the individual coordinates of the $d\times d$ symmetric Hessian matrix of $p_{xy}(\bm{\theta})$, that is,
$\frac{\partial^{2}p_{xy}(\bm{\theta})}{\partial{\bm{\theta}}\partial{\bm{\theta}}^{\top}}=\begin{pmatrix}\frac{\partial^{2}p_{xy}(\bm{\theta})}{\partial{\rho}^{2}}&\frac{\partial^{2}p_{xy}(\bm{\theta})}{\partial{\rho}\partial{a_{1}}}&\cdots&\frac{\partial^{2}p_{xy}(\bm{\theta})}{\partial{\rho}\partial{a_{K_{X}-1}}}&\frac{\partial^{2}p_{xy}(\bm{\theta})}{\partial{\rho}\partial{b_{1}}}&\cdots&\frac{\partial^{2}p_{xy}(\bm{\theta})}{\partial{\rho}\partial{b_{K_{Y}-1}}}\\ \frac{\partial^{2}p_{xy}(\bm{\theta})}{\partial{a_{1}}\partial{\rho}}&\frac{\partial^{2}p_{xy}(\bm{\theta})}{\partial{a_{1}}^{2}}&\cdots&\frac{\partial^{2}p_{xy}(\bm{\theta})}{\partial{a_{1}}\partial{a_{K_{X}-1}}}&\frac{\partial^{2}p_{xy}(\bm{\theta})}{\partial{a_{1}}\partial{b_{1}}}&\cdots&\frac{\partial^{2}p_{xy}(\bm{\theta})}{\partial{a_{1}}\partial{b_{K_{Y}-1}}}\\ \vdots&\vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ \frac{\partial^{2}p_{xy}(\bm{\theta})}{\partial{a_{K_{X}-1}}\partial{\rho}}&\frac{\partial^{2}p_{xy}(\bm{\theta})}{\partial{a_{K_{X}-1}}\partial{a_{1}}}&\cdots&\frac{\partial^{2}p_{xy}(\bm{\theta})}{\partial{a_{K_{X}-1}}^{2}}&\frac{\partial^{2}p_{xy}(\bm{\theta})}{\partial{a_{K_{X}-1}}\partial{b_{1}}}&\cdots&\frac{\partial^{2}p_{xy}(\bm{\theta})}{\partial{a_{K_{X}-1}}\partial{b_{K_{Y}-1}}}\\ \frac{\partial^{2}p_{xy}(\bm{\theta})}{\partial{b_{1}}\partial{\rho}}&\frac{\partial^{2}p_{xy}(\bm{\theta})}{\partial{b_{1}}\partial{a_{1}}}&\cdots&\frac{\partial^{2}p_{xy}(\bm{\theta})}{\partial{b_{1}}\partial{a_{K_{X}-1}}}&\frac{\partial^{2}p_{xy}(\bm{\theta})}{\partial{b_{1}}^{2}}&\cdots&\frac{\partial^{2}p_{xy}(\bm{\theta})}{\partial{b_{1}}\partial{b_{K_{Y}-1}}}\\ \vdots&\vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ \frac{\partial^{2}p_{xy}(\bm{\theta})}{\partial{b_{K_{Y}-1}}\partial{\rho}}&\frac{\partial^{2}p_{xy}(\bm{\theta})}{\partial{b_{K_{Y}-1}}\partial{a_{1}}}&\cdots&\frac{\partial^{2}p_{xy}(\bm{\theta})}{\partial{b_{K_{Y}-1}}\partial{a_{K_{X}-1}}}&\frac{\partial^{2}p_{xy}(\bm{\theta})}{\partial{b_{K_{Y}-1}}\partial{b_{1}}}&\cdots&\frac{\partial^{2}p_{xy}(\bm{\theta})}{\partial{b_{K_{Y}-1}}^{2}}\end{pmatrix}.$
This Hessian matrix can alternatively be expressed as follows, which follows by (A.2):
$\begin{split}&\frac{\partial^{2}p_{xy}(\bm{\theta})}{\partial{\bm{\theta}}\partial{\bm{\theta}}^{\top}}=\\ &\quad\frac{\partial^{2}}{\partial{\bm{\theta}}\partial{\bm{\theta}}^{\top}}\Phi_{2}\left(a_{x},b_{y};\rho\right)-\frac{\partial^{2}}{\partial{\bm{\theta}}\partial{\bm{\theta}}^{\top}}\Phi_{2}\left(a_{x-1},b_{y};\rho\right)-\frac{\partial^{2}}{\partial{\bm{\theta}}\partial{\bm{\theta}}^{\top}}\Phi_{2}\left(a_{x},b_{y-1};\rho\right)+\frac{\partial^{2}}{\partial{\bm{\theta}}\partial{\bm{\theta}}^{\top}}\Phi_{2}\left(a_{x-1},b_{y-1};\rho\right).\end{split}$ (A.3)
First, by means of repeated applications of the product rule and chain rule it can be shown that for any $u,v\in\mathbb{R}$,
$\frac{\partial^{2}}{\partial{\rho}^{2}}\Phi_{2}\left(u,v;\rho\right)=\frac{\partial}{\partial\rho}\phi_{2}\left(u,v;\rho\right)=\frac{\phi_{2}\left(u,v;\rho\right)}{(1-\rho^{2})^{2}}\Big{(}(1-\rho^{2})(\rho+uv)-\rho\big{(}u^{2}-2\rho uv+v^{2}\big{)}\Big{)},$
as well as
$\frac{\partial^{2}}{\partial{u}^{2}}\Phi_{2}\left(u,v;\rho\right)=\phi_{1}^{\prime}(u)\Phi_{1}\left(\frac{v-\rho u}{\sqrt{1-\rho^{2}}}\right)-\frac{\rho}{\sqrt{1-\rho^{2}}}\phi_{1}\left(u\right)\phi_{1}\left(\frac{v-\rho u}{\sqrt{1-\rho^{2}}}\right),$
where
$\phi_{1}^{\prime}(u)=-\frac{u}{\sqrt{2\pi}}\exp\left(-u^{2}/2\right),$
which follows immediately by the chain rule.
Next, for the second order cross-derivatives, it can be shown that
$\frac{\partial^{2}}{\partial{u}\partial{\rho}}\Phi_{2}\left(u,v;\rho\right)=\phi_{1}\left(u\right)\phi_{1}\left(\frac{v-\rho u}{\sqrt{1-\rho^{2}}}\right)\frac{\rho v-u}{(1-\rho^{2})^{3/2}}$
and
$\frac{\partial^{2}}{\partial{u}\partial{v}}\Phi_{2}\left(u,v;\rho\right)=\phi_{1}\left(u\right)\phi_{1}\left(\frac{v-\rho u}{\sqrt{1-\rho^{2}}}\right)\frac{1}{\sqrt{1-\rho^{2}}},$
both by applications of the chain rule and product rule.
It now follows by (A.3) combined with these second order cross-derivatives that
$\frac{\partial^{2}p_{xy}(\bm{\theta})}{\partial{a_{k}}\partial{\rho}}=\begin{dcases}\frac{\partial^{2}}{\partial{a_{x}}\partial{\rho}}\Phi_{2}\left(a_{x},b_{y};\rho\right)-\frac{\partial^{2}}{\partial{a_{x}}\partial{\rho}}\Phi_{2}\left(a_{x},b_{y-1};\rho\right)&\text{if }k=x,\\ -\frac{\partial^{2}}{\partial{a_{x-1}}\partial{\rho}}\Phi_{2}\left(a_{x-1},b_{y};\rho\right)+\frac{\partial^{2}}{\partial{a_{x-1}}\partial{\rho}}\Phi_{2}\left(a_{x-1},b_{y-1};\rho\right)&\text{if }k=x-1,\\ 0&\text{otherwise},\end{dcases}$
and
$\frac{\partial^{2}p_{xy}(\bm{\theta})}{\partial{a_{k}}\partial{b_{l}}}=\begin{dcases}\frac{\partial^{2}}{\partial{a_{k}}\partial{b_{l}}}\Phi_{2}\left(a_{x},b_{y};\rho\right)&\text{if }(k,l)\in\big{\{}(x,y),(x-1,y-1)\big{\}},\\ -\frac{\partial^{2}}{\partial{a_{k}}\partial{b_{l}}}\Phi_{2}\left(a_{x},b_{y};\rho\right)&\text{if }(k,l)\in\big{\{}(x-1,y),(x,y-1)\big{\}},\\ 0&\text{otherwise},\end{dcases}$
and
$\frac{\partial^{2}p_{xy}(\bm{\theta})}{\partial{a_{k}}\partial{a_{l}}}=\begin{dcases}\frac{\partial^{2}p_{xy}(\bm{\theta})}{\partial{a_{k}}^{2}}&\text{if }k=l,\\ 0&\text{otherwise},\end{dcases}$
and
$\frac{\partial^{2}p_{xy}(\bm{\theta})}{\partial{a_{k}}^{2}}=\begin{dcases}\frac{\partial^{2}}{\partial{a_{x}}^{2}}\Phi_{2}\left(a_{x},b_{y};\rho\right)-\frac{\partial^{2}}{\partial{a_{x}}^{2}}\Phi_{2}\left(a_{x},b_{y-1};\rho\right)&\text{if }k=x,\\ -\frac{\partial^{2}}{\partial{a_{x-1}}^{2}}\Phi_{2}\left(a_{x-1},b_{y};\rho\right)+\frac{\partial^{2}}{\partial{a_{x-1}}^{2}}\Phi_{2}\left(a_{x-1},b_{y-1};\rho\right)&\text{if }k=x-1,\\ 0&\text{otherwise},\end{dcases}$
and, finally,
$\frac{\partial^{2}p_{xy}(\bm{\theta})}{\partial{\rho}^{2}}=$
$\quad\frac{\partial^{2}}{\partial{\rho}^{2}}\Phi_{2}\left(a_{x},b_{y};\rho\right)-\frac{\partial^{2}}{\partial{\rho}^{2}}\Phi_{2}\left(a_{x-1},b_{y};\rho\right)-\frac{\partial^{2}}{\partial{\rho}^{2}}\Phi_{2}\left(a_{x},b_{y-1};\rho\right)+\frac{\partial^{2}}{\partial{\rho}^{2}}\Phi_{2}\left(a_{x-1},b_{y-1};\rho\right).$
Appendix B Remark on numerical optimization
[…]
Appendix C Additional results
Figure C.1 visualizes the mean squared error of all estimated parameters in the polychoric model in the first simulation in Section 6. We do not plot the results of the Pearson sample correlation coefficient because it does not estimate threshold parameters.
The remainder of this section contains additional results of the empirical application from Section 7. Table C.1 lists the unipolar markers of the three Big Five scales used by Arias et al., (2020), namely extroversion, conscientiousness, and neuroticism. Tables C.2C.4 contain the (polychoric) correlation matrices of the items in each scale, estimated by maximum likelihood and our robust estimator, while Figures C.2C.4 visualize the absolute difference between the two estimators for each pairwise correlation.
Figure C.1: Mean squared error of all estimated parameters in the polychoric model, $\|\widehat{\bm{\theta}}_{N}-\bm{\theta}_{*}\|^{2}$, in the first simulation in Section 6.
Adjective marker pairs
Construct Index Positive (P) Negative (N)
Extroversion (E) 1 extraverted introverted
2 energetic unenergetic
3 talkative silent
4 bold timid
5 assertive unassertive
6 adventurous unadventurous
Conscientiousness (C) 1 organized disorganized
2 responsible irresponsible
3 conscientious negligent
4 practical impractical
5 thorough careless
6 hardworking lazy
Neuroticism (N) 1 calm angry
2 relaxed tense
3 at ease nervous
4 not envious envious
5 stable unstable
6 contented discontented
Table C.1: Unipolar markers of three Big Five personality traits (Goldberg,, 1992). Each trait is measured by six pairs of items, where each item is a single English adjective. Each item pair consists of a positive and negative item. We explain item identifiers by means of the following example. Item “C3_P” refers to the positive (P) item in the 3rd pair of the conscientiousness (C) scale, that is, adjective “conscientious”, whereas “N1_N” would refer to “angry”.
N1_P N1_N N2_P N2_N N3_P N3_N N4_P N4_N N5_P N5_N N6_P N6_N
N1_P 1.00 -0.37 0.71 -0.50 0.69 -0.49 0.27 -0.24 0.58 -0.47 0.42 -0.32
N1_N -0.37 1.00 -0.40 0.55 -0.39 0.47 -0.19 0.40 -0.39 0.60 -0.32 0.56
N2_P 0.71 -0.40 1.00 -0.55 0.75 -0.54 0.26 -0.26 0.55 -0.41 0.53 -0.47
N2_N -0.50 0.55 -0.55 1.00 -0.54 0.65 -0.24 0.42 -0.41 0.57 -0.31 0.52
N3_P 0.69 -0.39 0.75 -0.54 1.00 -0.53 0.29 -0.28 0.63 -0.44 0.52 -0.48
N3_N -0.49 0.47 -0.54 0.65 -0.53 1.00 -0.28 0.43 -0.44 0.58 -0.29 0.47
N4_P 0.27 -0.19 0.26 -0.24 0.29 -0.28 1.00 -0.61 0.26 -0.20 0.18 -0.20
N4_N -0.24 0.40 -0.26 0.42 -0.28 0.43 -0.61 1.00 -0.33 0.46 -0.22 0.44
N5_P 0.58 -0.39 0.55 -0.41 0.63 -0.44 0.26 -0.33 1.00 -0.69 0.53 -0.46
N5_N -0.47 0.60 -0.41 0.57 -0.44 0.58 -0.20 0.46 -0.69 1.00 -0.35 0.57
N6_P 0.42 -0.32 0.53 -0.31 0.52 -0.29 0.18 -0.22 0.53 -0.35 1.00 -0.58
N6_N -0.32 0.56 -0.47 0.52 -0.48 0.47 -0.20 0.44 -0.46 0.57 -0.58 1.00
(a) Maximum likelihood estimates
N1_P N1_N N2_P N2_N N3_P N3_N N4_P N4_N N5_P N5_N N6_P N6_N
N1_P 1.00 -0.47 0.80 -0.58 0.79 -0.56 0.30 -0.26 0.63 -0.54 0.49 -0.39
N1_N -0.47 1.00 -0.48 0.58 -0.49 0.54 -0.26 0.45 -0.47 0.68 -0.43 0.63
N2_P 0.80 -0.48 1.00 -0.66 0.85 -0.60 0.32 -0.32 0.64 -0.50 0.60 -0.56
N2_N -0.58 0.58 -0.66 1.00 -0.70 0.76 -0.37 0.49 -0.48 0.60 -0.35 0.55
N3_P 0.79 -0.49 0.85 -0.70 1.00 -0.62 0.35 -0.39 0.66 -0.52 0.59 -0.57
N3_N -0.56 0.54 -0.60 0.76 -0.62 1.00 -0.42 0.49 -0.52 0.58 -0.37 0.53
N4_P 0.30 -0.26 0.32 -0.37 0.35 -0.42 1.00 -0.92 0.35 -0.30 0.30 -0.33
N4_N -0.26 0.45 -0.32 0.49 -0.39 0.49 -0.92 1.00 -0.39 0.50 -0.33 0.53
N5_P 0.63 -0.47 0.64 -0.48 0.66 -0.52 0.35 -0.39 1.00 -0.82 0.59 -0.55
N5_N -0.54 0.68 -0.50 0.60 -0.52 0.58 -0.30 0.50 -0.82 1.00 -0.44 0.61
N6_P 0.49 -0.43 0.60 -0.35 0.59 -0.37 0.30 -0.33 0.59 -0.44 1.00 -0.75
N6_N -0.39 0.63 -0.56 0.55 -0.57 0.53 -0.33 0.53 -0.55 0.61 -0.75 1.00
(b) Robust estimates
Table C.2: Estimated correlation matrices of the items in the neuroticism scale from the data in Arias et al., (2020, Sample 1; $N=725$) using MLE (top panel) and our robust estimator with tuning constant $c=1.6$ (bottom panel). The items are “calm” (N1_P), “angry” (N1_N), “relaxed” (N2_P), “tense” (N2_N), “at ease” (N3_P), “nervous” (N3_N), “not envious” (N4_P), “envious” (N4_N), “stable” (N5_P), “unstable” (N5_N), “contented” (N6_P), and “discontented” (N6_N). For the item naming given in parentheses, items with identical identifier (the integer after the first “N”) are polar opposites, where a last character “P” refers to the positive opposite and “N” to the negative opposite.
E1_P E1_N E2_P E2_N E3_P E3_N E4_P E4_N E5_P E5_N E6_P E6_N
E1_P 1.00 -0.77 0.50 -0.26 0.70 -0.50 0.56 -0.42 0.51 -0.40 0.51 -0.32
E1_N -0.77 1.00 -0.38 0.34 -0.59 0.61 -0.45 0.54 -0.47 0.50 -0.35 0.37
E2_P 0.50 -0.38 1.00 -0.65 0.43 -0.27 0.49 -0.28 0.47 -0.38 0.54 -0.39
E2_N -0.26 0.34 -0.65 1.00 -0.24 0.34 -0.30 0.40 -0.32 0.48 -0.38 0.49
E3_P 0.70 -0.59 0.43 -0.24 1.00 -0.59 0.44 -0.36 0.46 -0.40 0.41 -0.25
E3_N -0.50 0.61 -0.27 0.34 -0.59 1.00 -0.27 0.56 -0.35 0.45 -0.24 0.37
E4_P 0.56 -0.45 0.49 -0.30 0.44 -0.27 1.00 -0.41 0.64 -0.49 0.54 -0.34
E4_N -0.42 0.54 -0.28 0.40 -0.36 0.56 -0.41 1.00 -0.49 0.60 -0.27 0.40
E5_P 0.51 -0.47 0.47 -0.32 0.46 -0.35 0.64 -0.49 1.00 -0.71 0.39 -0.23
E5_N -0.40 0.50 -0.38 0.48 -0.40 0.45 -0.49 0.60 -0.71 1.00 -0.34 0.45
E6_P 0.51 -0.35 0.54 -0.38 0.41 -0.24 0.54 -0.27 0.39 -0.34 1.00 -0.68
E6_N -0.32 0.37 -0.39 0.49 -0.25 0.37 -0.34 0.40 -0.23 0.45 -0.68 1.00
(a) Maximum likelihood estimates
E1_P E1_N E2_P E2_N E3_P E3_N E4_P E4_N E5_P E5_N E6_P E6_N
E1_P 1.00 -0.87 0.55 -0.34 0.75 -0.62 0.58 -0.58 0.54 -0.45 0.55 -0.39
E1_N -0.87 1.00 -0.40 0.36 -0.67 0.63 -0.52 0.62 -0.52 0.51 -0.36 0.37
E2_P 0.55 -0.40 1.00 -0.84 0.50 -0.32 0.56 -0.38 0.55 -0.43 0.57 -0.44
E2_N -0.34 0.36 -0.84 1.00 -0.30 0.35 -0.40 0.43 -0.41 0.54 -0.45 0.53
E3_P 0.75 -0.67 0.50 -0.30 1.00 -0.71 0.50 -0.51 0.52 -0.50 0.42 -0.28
E3_N -0.62 0.63 -0.32 0.35 -0.71 1.00 -0.38 0.62 -0.47 0.47 -0.30 0.37
E4_P 0.58 -0.52 0.56 -0.40 0.50 -0.38 1.00 -0.55 0.73 -0.64 0.61 -0.48
E4_N -0.58 0.62 -0.38 0.43 -0.51 0.62 -0.55 1.00 -0.61 0.66 -0.33 0.44
E5_P 0.54 -0.52 0.55 -0.41 0.52 -0.47 0.73 -0.61 1.00 -0.85 0.44 -0.29
E5_N -0.45 0.51 -0.43 0.54 -0.50 0.47 -0.64 0.66 -0.85 1.00 -0.41 0.47
E6_P 0.55 -0.36 0.57 -0.45 0.42 -0.30 0.61 -0.33 0.44 -0.41 1.00 -0.83
E6_N -0.39 0.37 -0.44 0.53 -0.28 0.37 -0.48 0.44 -0.29 0.47 -0.83 1.00
(b) Robust estimates
Table C.3: Estimated correlation matrices of the items in the extroversion scale from the data in Arias et al., (2020, Sample 1; $N=725$) using MLE (top panel) and our robust estimator with tuning constant $c=1.6$ (bottom panel). The items are “extraverted” (E1_P), “introverted” (E1_N), “energetic” (E2_P), “unenergetic” (E2_N), “talkative” (E3_P), “silent” (E3_N), “bold” (E4_P), “timid” (E4_N), “assertive” (E5_P), “unassertive” (E5_N), “adventurous” (E6_P), and “unadventurous” (E6_N). For the item naming given in parentheses, items with identical identifier (the integer after the first “N”) are polar opposites, where a last character “P” refers to the positive opposite and “N” to the negative opposite.
C1_P C1_N C2_P C2_N C3_P C3_N C4_P C4_N C5_P C5_N C6_P C6_N
C1_P 1.00 -0.76 0.56 -0.42 0.34 -0.35 0.37 -0.26 0.51 -0.40 0.43 -0.43
C1_N -0.76 1.00 -0.51 0.58 -0.24 0.55 -0.32 0.44 -0.43 0.61 -0.44 0.55
C2_P 0.56 -0.51 1.00 -0.69 0.42 -0.55 0.57 -0.42 0.51 -0.54 0.65 -0.55
C2_N -0.42 0.58 -0.69 1.00 -0.39 0.75 -0.48 0.67 -0.42 0.70 -0.53 0.63
C3_P 0.34 -0.24 0.42 -0.39 1.00 -0.32 0.39 -0.34 0.44 -0.33 0.38 -0.25
C3_N -0.35 0.55 -0.55 0.75 -0.32 1.00 -0.37 0.59 -0.38 0.71 -0.44 0.53
C4_P 0.37 -0.32 0.57 -0.48 0.39 -0.37 1.00 -0.51 0.36 -0.38 0.39 -0.31
C4_N -0.26 0.44 -0.42 0.67 -0.34 0.59 -0.51 1.00 -0.38 0.59 -0.31 0.43
C5_P 0.51 -0.43 0.51 -0.42 0.44 -0.38 0.36 -0.38 1.00 -0.43 0.54 -0.39
C5_N -0.40 0.61 -0.54 0.70 -0.33 0.71 -0.38 0.59 -0.43 1.00 -0.43 0.53
C6_P 0.43 -0.44 0.65 -0.53 0.38 -0.44 0.39 -0.31 0.54 -0.43 1.00 -0.61
C6_N -0.43 0.55 -0.55 0.63 -0.25 0.53 -0.31 0.43 -0.39 0.53 -0.61 1.00
(a) Maximum likelihood estimates
C1_P C1_N C2_P C2_N C3_P C3_N C4_P C4_N C5_P C5_N C6_P C6_N
C1_P 1.00 -0.89 0.57 -0.56 0.36 -0.46 0.43 -0.35 0.54 -0.56 0.49 -0.52
C1_N -0.89 1.00 -0.58 0.64 -0.31 0.60 -0.38 0.47 -0.48 0.69 -0.52 0.61
C2_P 0.57 -0.58 1.00 -0.86 0.45 -0.68 0.62 -0.54 0.55 -0.65 0.69 -0.64
C2_N -0.56 0.64 -0.86 1.00 -0.44 0.80 -0.57 0.74 -0.50 0.76 -0.61 0.66
C3_P 0.36 -0.31 0.45 -0.44 1.00 -0.43 0.42 -0.46 0.17 -0.41 0.40 -0.26
C3_N -0.46 0.60 -0.68 0.80 -0.43 1.00 -0.48 0.70 -0.52 0.78 -0.55 0.59
C4_P 0.43 -0.38 0.62 -0.57 0.42 -0.48 1.00 -0.68 0.39 -0.47 0.44 -0.33
C4_N -0.35 0.47 -0.54 0.74 -0.46 0.70 -0.68 1.00 -0.47 0.66 -0.42 0.45
C5_P 0.54 -0.48 0.55 -0.50 0.17 -0.52 0.39 -0.47 1.00 -0.54 0.60 -0.45
C5_N -0.56 0.69 -0.65 0.76 -0.41 0.78 -0.47 0.66 -0.54 1.00 -0.59 0.61
C6_P 0.49 -0.52 0.69 -0.61 0.40 -0.55 0.44 -0.42 0.60 -0.59 1.00 -0.69
C6_N -0.52 0.61 -0.64 0.66 -0.26 0.59 -0.33 0.45 -0.45 0.61 -0.69 1.00
(b) Robust estimates
Table C.4: Estimated correlation matrices of the items in the conscientiousness scale from the data in Arias et al., (2020, Sample 1; $N=725$) using MLE (top panel) and our robust estimator with tuning constant $c=1.6$ (bottom panel). The items are “calm” (C1_P), “angry” (C1_N), “relaxed” (C2_P), “tense” (C2_N), “at ease” (C3_P), “nervous” (C3_N), “not envious” (C4_P), “envious” (C4_N), “stable” (C5_P), “unstable” (C5_N), “contented” (C6_P), and “discontented” (C6_N). For the item naming given in parentheses, items with identical identifier (the integer after the first “N”) are polar opposites, where a last character “P” refers to the positive opposite and “N” to the negative opposite.
Figure C.2: Difference between absolute estimates for the polychoric correlation coefficient of our robust estimator and the MLE for each item pair in the neuroticism scale, using the data of Arias et al., (2020). The items are “calm” (N1_P), “angry” (N1_N), “relaxed” (N2_P), “tense” (N2_N), “at ease” (N3_P), “nervous” (N3_N), “not envious” (N4_P), “envious” (N4_N), “stable” (N5_P), “unstable” (N5_N), “contented” (N6_P), and “discontented” (N6_N). For the item naming given in parentheses, items with identical identifier (the integer after the first “N”) are polar opposites, where a last character “P” refers to the positive opposite and “N” to the negative opposite.
Figure C.3: Difference between absolute estimates for the polychoric correlation coefficient of our robust estimator and the MLE for each item pair in the extroversion scale, using the data of Arias et al., (2020). The items are “extraverted” (E1_P), “introverted” (E1_N), “energetic” (E2_P), “unenergetic” (E2_N), “talkative” (E3_P), “silent” (E3_N), “bold” (E4_P), “timid” (E4_N), “assertive” (E5_P), “unassertive” (E5_N), “adventurous” (E6_P), and “unadventurous” (E6_N). For the item naming given in parentheses, items with identical identifier (the integer after the first “N”) are polar opposites, where a last character “P” refers to the positive opposite and “N” to the negative opposite.
Figure C.4: Difference between absolute estimates for the polychoric correlation coefficient of our robust estimator and the MLE for each item pair in the conscientiousness scale, using the data of Arias et al., (2020). The items are “calm” (C1_P), “angry” (C1_N), “relaxed” (C2_P), “tense” (C2_N), “at ease” (C3_P), “nervous” (C3_N), “not envious” (C4_P), “envious” (C4_N), “stable” (C5_P), “unstable” (C5_N), “contented” (C6_P), and “discontented” (C6_N). For the item naming given in parentheses, items with identical identifier (the integer after the first “N”) are polar opposites, where a last character “P” refers to the positive opposite and “N” to the negative opposite.
$X$\$Y$ 1 2 3 4 5
1 9,814,457,557.73 16,011.33 11.82 1.14 0.65
2 2,424.07 10.07 0.80 0.90 1.42
3 5.48 0.65 0.99 0.80 77.14
4 0.88 0.92 0.61 12.66 222,528.08
5 0.89 0.88 36.01 55,420.33 995,017,243,197.60
Table C.5: Decentralized Pearson residual, $\widehat{f}_{N}(x,y)\big{/}p_{xy}(\widehat{\bm{\theta}}_{N})$, of each response $(x,y)$ for the “not envious” ($X$)–“envious” ($Y$) item pair in the measurements of Arias et al., (2020) of a neuroticism scale. Estimate $\widehat{\bm{\theta}}_{N}$ was computed with tuning constant $c=1.6$.
Figure C.5: Dot plot of cells for the neuroticism item adjective pair “envious” and “not envious” in the data of Arias et al., (2020), where each item has five Likert-type response options, anchored by “very inaccurate” (= 1) and “very accurate” (= 5). Each dot’s size is proportional to the relative empirical frequency of its associated contingency table cell, $\widehat{f}_{N}(x,y)$, whereas its color varies by the value of the cell’s DPR, $\widehat{f}_{N}(x,y)/p_{xy}(\widehat{\bm{\theta}}_{N})$, at robust parameter estimate with tuning constant $c=1.6$. Note that some cells have an extremely large DPR (substantially larger than ideal value 1; see also Table C.5), indicating a poor model fit for those cells. We therefore censored the color scale so that a fully saturated red is assigned to all cells with DPR $\geq 4$, i.e., the empirical frequency $\widehat{f}_{N}(x,y)$ being at least four times as large as the model frequency $p_{xy}(\widehat{\bm{\theta}}_{N})$.